### v5

# Large Language Models Are Struggle to Cope with Unreasonability in Math Problems

Abstract
Recent research have demonstrated LLMs’ impressive performance in math and reasoning. However, the capacity of LLMs to address math problems under unconventional conditions, such as internal inconsistencies and flawed assumptions, remains largely unexplored. In this paper, we propose a novel benchmark Unreasonable Math Problem (UMP) designed to assess LLMs’ ability to recognize and respond to unreasonability in math problem. The benchmark consists of a carefully curated collection of unreasonable math questions across diverse types. Based on extensive experiments covering 19 LLMs, we observe that even state-of-the-art models such as GPT-4o achieve only limited performance of 0.6 in UMP, while reasoning models such as DeepSeek-R1 are prone to overthinking and unstable. We further explore strategies for improving the recognition of unreasonable inputs, shedding light on both the possibility and limitations of LLMs in this challenging setting.
Large Language Models Are Struggle to Cope with Unreasonability
in Math Problems
Jingyuan Ma1 , Damai Dai1 , Zihang Yuan2, Rui Li1, Weilin Luo3, Bin Wang3, Qun Liu3, Lei Sha2, Zhifang Sui1 1National Key Laboratory for Multimedia Information Processing, School of Computer Science, Peking University 2Institute of Artificial Intelligence, Beihang University 3Huawei Noah’s Ark Lab, China {mjy@stu.pku.edu.cn}
1 Introduction
Large language models (LLMs) have recently achieved remarkable progress in math and reasoning, particularly with the emergence of reasoning-enhanced models like O1 and R1 DeepSeek-AI (2025), demonstrate increasingly structured and deliberate reasoning capabilities. As a result, current LLMs have surpassed earlier baselines and achieved over 90% accuracy on math benchmarks like GSM8K Cobbe et al. (2021a) and MATH Hendrycks et al. (2021).
However, despite this impressive performance on challenging math problems, we find that these models often struggle to recognizing when a math question is unreasonable. Even when the irrationality, such as a logical contradiction or a violation of basic constraints, is immediately obvious to humans. LLMs frequently follow the misleading question setup, producing confident yet nonsensical answers, or becoming trapped in cycles of excessive reasoning that fail to converge. This counter-intuitive behavior raises serious concerns about the reliability in their practical applications such as automated tutoring Kasneci et al. (2023), early education Zhang et al. (2024b), or open-domain problem solving Lin and Chen (2023).
To enable a comprehensive analysis of how LLMs behave when confronted with mathematically unreasonable inputs, we introduce the Unreasonable Math Problems (UMP) benchmark. We take questions from GSM8K Cobbe et al. (2021a), MATH Hendrycks et al. (2021), AIME24 MAA (2024), and AMC23 to generated unreasonable variants. Specifically, we utilize GPT-4o to apply minimal edits that introduce specific forms of unreasonability while preserving the original questions structure. These edits are guided by a set of rule-based transformation templates. All generated questions are verified to ensure that they are genuinely unreasonable in a mathematical or logical sense, and that they remain similar to the original questions in length, topic, and surface form. This careful construction allows us to attribute differences in model behavior specifically to the presence of unreasonableness, rather than to confounding changes in difficulty.
Based on UMP benchmark, we conducted extensive experiments on various models spanning three categories: general-purpose models (e.g., GPT-4oOpe (2023)), reasoning-enhanced models (e.g., DeepSeek-R1DeepSeek-AI (2025)), and math-specialized models (e.g., Qwen-MathYang et al. (2024a)). In addition, we analyze LLMs’ failure patterns according to tokens repetition, reflection frequency, and token entropy. We find: 1) general-purpose models often proceed confidently without recognizing the unreasonableness of the question; 2) reasoning-enhanced models tend to overthink and fall into excessive self-correction; 3) math-specialized models may fail to initiate reasoning when confronted with unreasonable premises. We further explored several mitigation strategies but found that none could robustly resolve these failure modes without introducing trade-offs, such as decreased performance on standard inputs.
Our main contributions are as follows:
We propose the Unreasonable Math Problems (UMP) benchmark by minimally editing standard math datasets to create diverse and comprehensive unreasonable problems.
We conduct comprehensive experiments and multi-dimensional analyses to examine how different types of LLMs behave when presented with unreasonable questions.
We identify distinct failure modes and reveal that models with strong performance on well-posed tasks often behave unstably or fail to recognize unreasonableness when assumptions are flawed.
We explore prompting and fine-tuning as strategies to investigate the potential of LLMs to detect unreasonableness.
2 Unreasonable Math Problems(UMP) Benchmark
While large language models have demonstrated strong performance on standard mathematical reasoning benchmarks, they often produce inaccurate or confusing responses when presented with mathematically unreasonable problems, questions that contain flawed assumptions, undefined variables, or logical inconsistencies (as show in Figure 1). To systematically evaluate model behavior under such conditions, we construct the Unreasonable Math Problems (UMP) benchmark, which consists of 1000+ unreasonable math problems, each paired with its corresponding original version, focusing on assessing LLMs’ ability to detect and respond to irrational inputs.
2.1 Data Category
We identify five prevalent types of mathematical unreasonableness commonly found in LLM failure cases: (1) undefined variables, (2) illogical scenarios, (3) incorrect assumptions, (4) misinterpretation of units, and (5) inconsistent conditions. Each instance in our benchmark is represented as a quintuple , consisting of the original question and its answer , the unreasonable variant with its assigned type , and an explanation describing why is unreasonable. Detailed definitions and examples for each category are provided in Appendix A.
2.1.1 Data Generation
Our data generation process is inspired by MetaMath (Yu et al., 2023b), which leverages LLMs to produce problem variants under controlled transformations. As illustrated in Figure 2, we begin with test set questions drawn from four widely-used math benchmarks: GSM8K Cobbe et al. (2021a), a collection of grade-school level problems; MATH Hendrycks et al. (2021), which covers formal secondary school mathematics; and AIME24MAA (2024) and AMC23, both of which contain high level competition problems with symbolic or abstract formats. We manually construct a set of transformation rules corresponding to five types of mathematical unreasonableness and use them to guide an LLM in generating unreasonable variants for each original question, along with natural language explanations of why the modified version is irrational. To ensure the unreasonable variant remains close in surface form to the original, we compute cosine similarity between the sentence embeddings of the original and generated questions using SimCSE (Gao et al., 2021), a contrastively trained BERT-based modelDevlin et al. (2019). Only variants whose similarity exceeds a predefined threshold are retained for human verification.
2.1.2 Data Checking
We perform manual validation with the assistance of three annotators, each with at least undergraduate-level mathematical training. For each example, annotators are provided with the original problem, its unreasonable variant, and a corresponding explanation, and are asked to verify whether the variant introduces genuine mathematical unreasonableness while preserving surface similarity. Only examples that meet both criteria are retained in the final benchmark. Our transformation rules are deliberately designed to introduce flaws that are logically problematic. As such, we filter out cases that either fail to introduce meaningful irrationality or make the unreasonableness too explicit, for example, variants that contain phrases like “-3 people” or “2.5 people,” which violate commonsense constraints in an obvious way. In these cases, we revise the questions by modifying entities, adjusting numerical values, or refining the phrasing to ensure that the unreasonable elements remain logically embedded but superficially plausible.
2.1.3 Data Composition
Table 1 summarizes the distribution of unreasonableness types across different source datasets in the UMP benchmark. A large proportion of examples fall into the categories of Incorrect Assumptions (IA) and Inconsistent Conditions (IC). This distribution emerges from the model generation and human filtering process. In particular, these two types of flaws tended to produce more plausible and contextually coherent questions, making them more likely to be retained during human recheck. In comparison, AMC23 and AIME24 contain relatively few test questions (with AIME24 having only 30), which naturally limits the number of examples that can be derived from these datasets. Moreover, problems in these sources are often abstract and symbolic, making them less amenable to realistic and controlled perturbation. Rather than artificially forcing unreasonable variants in such cases, we adopt a conservative strategy and include only those examples where natural modifications can be applied without compromising the original structure. This selective inclusion ensures that the final benchmark maintains both diversity and quality across a range of mathematical domains.
| Dataset | UV | IS | IA | MU | IC | Total |
| GSM8K | 14.8% | 7.3% | 36.8% | 16.7% | 24.3% | 682 |
| MATH | 7.2% | 4.4% | 41.7% | 13.1% | 33.6% | 405 |
| AMC23 | 0.0% | 4.3% | 43.5% | 4.3% | 47.8% | 23 |
| AIME24 | 10.5% | 0.0% | 31.6% | 0.0% | 57.9% | 19 |
| Total | 8.1% | 4.0% | 38.4% | 8.5% | 40.9% | 1129 |
3 Experiments
3.1 Experimental Setup
We evaluate a diverse set of LLMs spanning three major categories. General-purpose models are primarily trained for instruction following and open-domain tasks; this group includes DeepSeek-V3 DeepSeek-AI (2024), GPT-4o Ope (2023), Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct Qwen Team (2024), and LLaMA3.1-8B-Instruct AI (2024). Math-specialized models are fine-tuned on mathematical corpora and optimized for numerical reasoning; we include Qwen2.5-Math-1.5B-Instruct, Qwen2.5-Math-7B-Instruct Yang et al. (2024a), and DeepSeek-Math-7B-Instruct Shao et al. (2024). Reasoning-enhanced models are designed to support complex multi-step reasoning and include DeepSeek-R1 and its distilled variants (e.g., DeepSeek-R1-Distill-Qwen-7B) DeepSeek-AI (2025), as well as QwQ-32B-Preview Team (2025) and Grok-3-Reasoning. All models are evaluated on both the original problems and their unreasonable variants.111Details of model versions and inference configurations are provided in Appendix C.
3.2 Evaluation Metrics
Evaluating model performance on unreasonable questions poses a fundamentally different challenge compared to standard accuracy-based evaluation. In conventional settings, the goal is to assess whether the model arrives at the correct answer. However, for unreasonable questions, the objective is to evaluate whether the model can recognize the irrationality embedded in the problem and respond appropriately by rejecting, questioning, or flagging the input as flawed. To address this, we introduce two complementary metrics:
Absolute score measures the proportion of unreasonable questions for which the model explicitly identifies the problem as flawed or unanswerable.
Relative score conditions on the model’s ability to correctly solve the original version of a question. It is defined as the proportion of corresponding unreasonable variants that are correctly recognized as unreasonable, among those for which the model answers the original (reasonable) version correctly. This design controls for a factor that in some cases, the irrationality in a problem only becomes apparent during intermediate reasoning steps. A model that lacks the necessary mathematical competence may never reach the point where the irrationality is revealed. By restricting the evaluation to questions the model can already solve in their original form, relative score isolates the model’s ability to detect unreasonableness from its general problem-solving skill.
| Model | GSM | MATH | AMC+AIME | AVG | ||||||||
| Acc | Abs | Rel | Acc | Abs | Rel | Acc | Abs | Rel | Acc | Abs | Rel | |
| General Models | ||||||||||||
| Qwen2.5-3B-Instruct | ||||||||||||
| Qwen2.5-7B-Instruct | ||||||||||||
| Llama-3.1-8B-Instruct | ||||||||||||
| DeepSeek-v3 | ||||||||||||
| GPT-4o | 0.610 | |||||||||||
| Claude-3.5-Sonnet | ||||||||||||
| Reasoning Models | ||||||||||||
| R1-Distill-Qwen-1.5B | ||||||||||||
| Marco-o1 | ||||||||||||
| R1-Distill-Qwen-7B | ||||||||||||
| R1-Distill-Llama-8B | ||||||||||||
| R1-Distill-Qwen-32B | ||||||||||||
| QwQ-32B-Preview | ||||||||||||
| DeepSeek-R1 | ||||||||||||
| Grok3-Reasoning | 0.903 | |||||||||||
| Math Models | ||||||||||||
| Qwen2.5-Math-1.5B-Instruct | 0.376 | |||||||||||
| MetaMath-Mistral-7B | ||||||||||||
| DeepSeek-Math-7B-Instruct | ||||||||||||
| NuminaMath-7B-CoT | ||||||||||||
| Qwen2.5-Math-7B-Instruct | ||||||||||||
3.2.1 Absolute Score
To evaluate how well a model identifies the unreasonableness of a given question, we adopt the LLM-as-a-judge paradigm Zheng et al. (2023), using DeepSeek-V3 to assess the quality of generated responses. Instead of a binary correct/incorrect judgment, we use a three-level rubric to better capture the nuances of model behavior:
A: Correctly identifies the main source of unreasonableness, provides a coherent explanation, and proposes a valid correction without introducing new errors.
B: Partially recognizes the unreasonableness or provides vague justification, but the explanation contains gaps, circular logic, or minor inconsistencies.
C: Fails to recognize the core flaw in the problem, misinterprets the reasoning, or introduces new contradictions in the explanation or correction.
Each model response receives a label , and we define a soft scoring function to assign partial credit:
The final absolute score is computed as the average soft score across all unreasonable questions:
To ensure alignment with our evaluation criteria, we design an in-context learning (ICL) prompt that includes annotated examples for each rating level, following best practices for LLM-based evaluation Dong et al. (2024).
3.2.2 Relative Score
While the absolute score provides a general measure of robustness, it does not account for the model’s underlying mathematical competence. In many cases, the flaw in an unreasonable question only becomes apparent during intermediate reasoning steps. If a model lacks the skill to solve the original problem, it may never reach the point where the flaw is revealed.
To control for this confound, we introduce the relative score, which measures a model’s ability to detect unreasonableness on the subset of questions it has already solved correctly in their original form. Formally, following Yang et al. (2024b), we define the relative score as:
Under a multi-sample evaluation setting, we generate responses for each original question and compute its average accuracy:
where indicates whether the -th sampled response to question is correct (1 if correct, 0 otherwise), and denotes the proportion of correct responses over samples.
We then define a set of confidently solved questions by applying a threshold to the average accuracy:
This ensures that only questions the model has solved reliably in their original (reasonable) form are considered.
Let be the set of unreasonable variants derived from question . Each variant is scored with a rating , which is mapped to a soft score via the function defined earlier.
The final relative score is computed by averaging the soft scores over all unreasonable variants associated with confidently solved questions:
3.3 Experimental Results
Table 2 reports model performance on UMP benchmark and three evaluation metrics: accuracy on original problems (Acc), and absolute and relative scores (Abs, Rel) on unreasonable variants. As defined in Section 3.2, these behavioral scores capture a model’s ability to detect and respond appropriately to flawed problem setups.
General-purpose models.
Models such as GPT-4o, Claude-3.5-Sonnet, and DeepSeek-v3 perform well on well-posed problems (e.g., GPT-4o: 0.863 Acc) and show moderate ability to detect unreasonableness (0.603 Abs). However, smaller models like Qwen2.5-3B and LLaMA-3.1-8B struggle with behavioral metrics, indicating difficulty in recognizing flawed logic. This suggests that robustness to irrationality is not a natural outcome of instruction tuning and may require additional improvement.
Reasoning-enhanced models.
These models achieve consistently high scores, with DeepSeek-R1 and grok-3-reasoning reaching 0.813 Abs and 0.903 Rel, respectively. Their structured reasoning and verification steps appear effective for detecting logical flaws, though often at the cost of longer, more verbose outputs (see Section 4.2).
Math-specialized models.
Despite strong accuracy on original problems (e.g., Qwen2.5-Math-7B: 0.894 Acc), this group generally underperforms on unreasonable problems(score mostly below 0.30), revealing a gap between reasoning proficiency and irrationality detection. MetaMath-Mistral-7B exemplifies this limitation, with only 0.139 on Rel.
Summary.
We observe a trade-off between task accuracy and robustness to unreasonableness. Reasoning-enhanced models excel at flaw detection, while math-specialized models, despite high precision on well-posed problems, often fail to recognize inconsistencies in adversarial variants.
4 Analysis
4.1 Does Solving the Original Help?
We investigate whether a model’s ability to solve the original version of a problem affects its capacity to detect flaws in the corresponding unreasonable variants. Intuitively, some flaws such as contradictions in intermediate steps may only be visible after progressing through the correct reasoning path. This motivates a direct comparison: how often does a model correctly identify unreasonableness when it has solved the original question, versus when it has not? To measure this, we revisit the set from Section 3.2, containing original questions that the model solves correctly with high confidence. We define its complement as the set of questions the model fails to solve. For each group, we compute the average score over their unreasonable variants, using the same scoring function . We then define the robustness gap as:
where and denote the average scores over and , respectively. A larger gap indicates that the model’s ability to detect unreasonableness depends more strongly on its success in solving the original problem.
4.2 Reasoning Models Tend to Overthink on Unreasonable Problems
A prominent behavioral failure we observe is overthinking, where the model enters repeated cycles of reflection and revision without making meaningful progress Sui et al. (2025); Cuadron et al. (2025). We quantify overthinking by counting occurrences of reflective phrases such as “rethink” or “I misunderstood,” which indicate mid-generation self-correction (see Appendix D for keyword details). Although we evaluate reflection frequency across all model types, only reasoning-enhanced models exhibit consistent and substantial increases on unreasonable problems.
As shown in Figure 3, reasoning models show an average increase in reflective behavior by a factor of three to four when moving from original to unreasonable questions. Instead of terminating early, these models tend to reevaluate their reasoning repeatedly, leading to longer and less coherent responses. This pattern is largely absent in general-purpose and math-specialized models, suggesting that stronger reasoning capacity also increases susceptibility to overthinking when faced with unreasonable inputs. 222Due to the limited number of AMC and AIME samples, we report detailed scores only for GSM and MATH datasets.
4.3 Lexical Collapse and Redundancy in Response
In addition to analyzing reflection frequency, we further examine how model outputs deteriorate under unreasonable conditions by introducing two complementary indicators: normalized token entropy and token-level repetition. These metrics are designed to quantify lexical diversity and redundancy, offering additional perspectives on generation instability across model types.
Normalized Token Entropy.
To assess lexical diversity, we compute normalized token entropy based on the empirical token frequency distribution in each response. For an output sequence of length , let be the count of token , and define its empirical probability as . Entropy is calculated as:
The normalization ensures comparability across varying output lengths. A lower indicates token concentration, which often signals collapsed or repetitive responses. Conversely, higher entropy suggests fluent and lexically diverse outputs Yuan et al. (2024).
As shown in Figure 4, most models exhibit a decline in entropy on unreasonable questions. This drop is especially pronounced in reasoning-enhanced models such as R1-Distill-Qwen-32B, where entropy decreases by 0.1–0.15 on average. These collapsed outputs often stem from repeated, ineffective reasoning loops, which reduce informativeness and mask the model’s failure to recognize flawed assumptions. Examples of such behavior are presented in Appendix I.1.
Token-Level Repetition.
We further quantify redundancy by computing the number of repeated -grams (with ) within each output. High repetition reflects local generation loops, where the model reiterates similar phrasing instead of progressing logically.
Figure 4 shows that almost all models exhibit increased 10-gram repetition under unreasonable inputs, with the most substantial jumps observed in reasoning-enhanced models. In some cases, repetition increases by a factor of 5 to 10 compared to original problems, reinforcing the pattern of collapsed outputs. Interestingly, Qwen2.5-Math-7B-Instruct, a math-specialized model with strong performance on well-posed questions, also shows a marked rise in repetition, suggesting that semantic instability is not limited to reasoning-heavy models. Appendix I.2 provides detailed case studies of this failure pattern.
5 Can LLMs Detect Irrationality?
Through the experiments above, we observe that most models perform poorly at recognizing unreasonableness in mathematical problems. To explore possible ways to improve this behavior, we begin with a simple probing experiment to test whether models possess the basic ability to judge flawed inputs. Based on this, we further investigate two strategies, prompt-based intervention and supervised fine-tuning—to enhance the model’s capacity to detect irrationality.
5.1 Probing Study
Before exploring how to improve model behavior on unreasonable problems, we first ask a fundamental question: do models even understand what counts as unreasonable? If a model lacks this awareness, it is unlikely to respond appropriately when encountering flawed questions. To test this, we design a simple probing task, where the model is presented with short, isolated expressions that are syntactically valid but semantically unreasonable—for example, “the voting result is -20 votes” or “there are 2.5 people.” We refer to this metric as Unreasonable Phrase Accuracy (UPA). For each expression, the model is asked to judge whether it makes sense.
We construct this test set by extracting unreasonable phrases from the UMP benchmark and evaluate a representative subset of models used in our main experiments (Section 3), including general-purpose, reasoning-enhanced, and math-specialized models. As shown in Figure 5, general-purpose and reasoning-enhanced models typically perform well on this task, suggesting that they retain basic commonsense priors. In contrast, several math-specialized models perform noticeably worse, indicating that domain-specific fine-tuning may come at the cost of general semantic awareness.
5.2 Prompting LLMs with a Critical Thinking Signal
To explore lightweight methods for improving flaw detection, we experiment with inserting a simple instruction—such as “Please solve these problems with criticism.”—into the prompt. This encourages the model to approach the task with a more skeptical and reflective reasoning style. We evaluate this strategy on Qwen2.5-7B-Instruct and R1-Qwen-7B, and observe a consistent improvement in identifying unreasonable questions. As shown in Figure 6, both models achieve significantly higher Absolute Scores on unreasonable inputs, while their accuracy on original questions remains largely unaffected. Notably, with the critical-thinking prompt, the general-purpose Qwen2.5-7B-Instruct nearly matches the performance of the much larger DeepSeek-Chat in detecting flawed assumptions—highlighting the potential of prompt-based interventions to activate latent reasoning ability without additional training.
5.3 Fine-Tuning Based Method
We further investigate whether the model’s ability to detect irrationality can be enhanced through supervised fine-tuning. To construct the training data, we pair unreasonable questions from GSM8K with carefully edited responses based on DeepSeek-R1 outputs. These responses explicitly identify the flaws through concise, step-by-step reasoning, encouraging appropriate reflection without triggering overthinking. We also include original questions paired with similarly edited answers to maintain general reasoning competence and prevent the model from overgeneralizing flaw detection to well-posed inputs. This balanced training setup provides the model with positive demonstrations of how to respond to both flawed and valid inputs. We fine-tune on a subset of GSM8K and evaluate the resulting model on MATH to assess generalization beyond the training distribution. As shown in Figure 6, the fine-tuned model exhibits clear improvement in detecting unreasonable inputs. However, this gain comes at a cost: accuracy on original problems drops noticeably. This suggests a trade-off between flaw sensitivity and general problem-solving ability. Further training configurations and examples are detailed in Appendix F.
Discussion.
We adopt several commonly used enhancement strategies, including prompting and fine-tuning, both of which yield promising gains in unreasonable problem detection. However, these improvements consistently come at the cost of reduced accuracy on original, well-posed inputs. This trade-off highlights a core challenge: simple prompting or fine-tuning alone is insufficient to robustly address unreasonable inputs without introducing new side effects. The limitations of these methods underscore the inherent difficulty of the UMP benchmark and point to the need for more principled, generalizable flaw detection mechanisms that can integrate critical reasoning while preserving performance on standard tasks.
6 Related Work
6.1 LLMs as Math Problem Solvers
Large language models have demonstrated impressive capabilities in solving math word problems and symbolic reasoning tasks. The most widely used benchmarks in this space include GSM8K (Cobbe et al., 2021b), which targets elementary school-level problems with step-by-step annotations, and MATH (Hendrycks et al., 2021), which evaluates high-school competition-level mathematics. To improve performance on these tasks, various data-centric approaches have been proposed. WizardMath (Luo et al., 2023a) and MetaMath (Yu et al., 2023b) leverage self-instruct and verifier-guided generation to create high-quality training examples. These augmentations expose models to diverse problem formats and encourage generalization. In parallel, a wave of domain-specialized models has emerged—such as Qwen-MathYang et al. (2024a), DeepSeek-MathShao et al. (2024), and WizardMath-v2Luo et al. (2023b), which are fine-tuned on large-scale mathematical corpora and incorporate structured reasoning or reflection mechanisms. These models outperform general-purpose LLMs on math benchmarks.
6.2 Improve Models’ Inference Ability
Recent efforts have focused on enhancing the reasoning capabilities of large language models by structuring and extending their inference chains. The chain-of-thought (CoT) prompting method (Wei et al., 2022) was among the earliest approaches to guide models toward step-by-step reasoning. Subsequent variants such as Complex CoT (Fu et al., 2023) and Plan-and-Solve (Wang et al., 2023) further emphasized intermediate planning and decompositional reasoning. More recently, large-scale models like O1/R1 (DeepSeek-AI, 2025) and QwQ-32B (Team, 2025) have shown that extending the reasoning trajectory through structured plans, iterative self-correction, or reflective feedback can significantly improve performance on complex tasks. These models often adopt long-form generation and multi-phase solving strategies, which mimic human deliberation. In addition, verifier-based strategies such as Outcome-Supervised Learning (Yu et al., 2023a) and post-hoc reflection mechanisms have been introduced to scrutinize intermediate steps and promote robust decision-making, especially when the model’s output chain may contain errors.
6.3 Model Refusal Phenomenon
The phenomenon of model refusal, where LLMs decline to answer certain prompts is central to ensuring AI safety and reliability. Refusals may stem from safety training, uncertainty or knowledge gaps or defects in the prompt itself. Recent studies have explored both linear and nonlinear mechanisms behind refusal behaviors, including "refusal directions" in activation space and the use of "refusal tokens" to control response behavior at inference. However, over-refusal—where benign prompts are incorrectly rejected remains a key challengePasch (2025); Hildebrandt et al. (2025). Approaches like RAIT, CRaFT, and Think-Before-Refusal aim to reduce false refusals by incorporating reasoning and certainty estimationZhu et al. (2024); Zhang et al. (2024a). Evaluation efforts have led to dedicated benchmarks such as OR-Bench, SORRY-Bench, and MM-UPD, which assess refusal quality under various conditions, highlighting the need for context-aware and nuanced assessment frameworksCui et al. (2024); Xie et al. (2025); Miyai et al. (2025).
7 Conclusion
In this work, we investigate how large language models behave when confronted with unreasonable mathematical problems. To facilitate this, we construct a benchmark comprising over 1,000 questions containing hidden irrationality. Experimental results reveal that most models struggle to identify such unreasonableness. While reasoning-enhanced models are more likely to uncover hidden inconsistencies during multi-step reasoning, they often fall into repetitive reflection and overthinking behavior that ultimately hinders clarity and usefulness. To further explore whether models possess latent capabilities for detecting flaws, we experiment with critical-thinking prompts and supervised fine-tuning. Our findings suggest that many models do have the potential to detect unreasonable content, but this ability requires explicit activation. We hope that our benchmark serves as a valuable tool for evaluating both the reliability and behavioral robustness of LLMs in the presence of unreasonable inputs.
Limitations
This study has several limitations. First, we do not include proprietary reasoning models such as those from OpenAI in our analysis of reasoning-enhanced systems. Since these APIs do not expose intermediate reasoning chains, comparing their behavior to open-source models that explicitly generate multi-step thoughts would introduce unfairness in both behavioral and metric-based evaluations. Second, due to hardware constraints, our intervention experiments such as prompt engineering and supervised fine-tuning are conducted only on 7B-scale models. While these experiments provide encouraging insights, scaling these methods to larger models and exploring more effective solutions for mitigating failure behaviors remain important directions for future work.
Ethical Considerations
Our paper explores how llms perform in the face of unreasonable mathematical problems. Sometimes the model may not realize the irrationality in the math problem, so it may produce wrong answers and mislead the user.
Appendix A Category of UMP Benchmark
Below is the description of the five categories. Each of our data pieces is made up of five parts. An original question paired with its corresponding answer , along with a designated type of unreasonableness we aim to investigate, we systematically construct a rephrased, unreasonable question and provide an accompanying explanation . Consequently, our benchmark comprises quintuples of the form .
Appendix B Probabilistic View from Prior Conditioning
We frame the model’s ability to recognize unreasonable problems as a conditional behavior dependent on its understanding of the original question. Let denote the correctness score for a model response to a variant , assigned by an LLM. To reason about this behavior probabilistically, we apply the law of total expectation:
This formulation decomposes the expected quality of model behavior into two conditional components:
The expectation over questions the model can solve (), and
The expectation over questions it fails to solve ().
Our focus is on the first term, , which reflects the robustness of the model under the prior belief that it understands the base question.
Since each original question may yield multiple unreasonable variants, we define an empirical estimator for this conditional expectation by averaging over the scores of all variants:
This is a two-level Monte Carlo approximation of the conditional expectation. It provides a graded measure of how reliably a model detects semantic flaws, assuming sufficient prior understanding of the original problem.
Appendix C Model Versions and Inference Settings
We include both API-access and open-source models in our evaluation. The API versions used are gpt-4o-2024-11-20, grok-3-reasoning and deepseek-r1-2025-01-20, claude-3.5-sonnet-20241022. To maximize accuracy on original (well-posed) questions, we adopt temperature = 0.0(greedy decoding) for all open-source non-reasoning models. However, for reasoning-enhanced models, we observe that greedy decoding may result in unstable behaviors, such as excessively long outputs, especially in unreasonable problems. Based on these findings, we set temperature = 0.6 for all reasoning models during evaluation.(We provide further analysis in Appendix H.) We set , which means that the model must get all three responses right. For API-based models, we maintain their default decoding settings. Due to access limitations and cost constraints, all the responses are generated in a single pass without sampling-based aggregation.
Appendix D Keywords for Reflection Detection
To quantify reflective behavior in model responses, we apply a keyword-based heuristic to detect phrases associated with self-correction, doubt, or reevaluation. The following list of keywords is used to flag reflection:
sorry, apologize, actually, wait, let me, i made, mistake, error, incorrect, again, redo, retry, revise, correct, fix, adjust, modify, amend, update, let’s try, recalculate, recompute, reassess, reevaluate, reanalyze, re-examine, i was wrong, i erred, i miscalculated, i misunderstood, i misread, i overlooked, rethink, reconsider, reflect, review, in fact, on second thought, upon reflection, second guess, double check, verify, confirm, clarify, perhaps, maybe, possibly, alternatively, otherwise, instead, rather, my bad, my fault, my error, my mistake, i stand corrected, i take it back
This keyword set was manually curated to capture a wide range of reflective behaviors and was applied in our analysis of model overthinking.
Appendix E Similarity between Original Problems and Unreasonable Problems
To ensure that our constructed unreasonable problems remain as close as possible to their original counterparts, we conducted a similarity analysis between each pair of , where is the original question and is its unreasonable variant. We use cosine similarity over sentence embeddings (computed using a pre-trained BERT encoder) as our similarity metric. Figure 7 shows the distribution of similarity scores across all problem pairs. Most unreasonable problems achieve high similarity scores with their original versions, with the majority falling in the range, which means a threshold . This confirms that the perturbations introduced are minimal in surface form, allowing us to isolate the effects of semantic difference while keeping the wording largely intact.
| Model | |||
| DS-LLaMA-8B | 0.593 | 0.514 | 0.079 |
| DS-Math-7B | 0.158 | 0.064 | 0.094 |
| QwQ-32B | 0.560 | 0.542 | 0.019 |
| Qwen2.5-7B | 0.440 | 0.433 | 0.007 |
| Marco-o1 | 0.546 | 0.481 | 0.065 |
| DS-Qwen-7B | 0.533 | 0.437 | 0.096 |
| DS-Qwen-1.5B | 0.397 | 0.310 | 0.087 |
| Qwen2.5-3B | 0.339 | 0.230 | 0.109 |
| DS-Qwen-32B | 0.734 | 0.798 | -0.064 |
| DS-Chat | 0.562 | 0.611 | -0.049 |
| Numina-7B | 0.320 | 0.204 | 0.116 |
| DS-Reasoner | 0.844 | 0.806 | 0.039 |
| MetaMath-7B | 0.146 | 0.086 | 0.060 |
| Qwen2.5-Math-7B | 0.231 | 0.135 | 0.096 |
| GPT-4o | 0.657 | 0.610 | 0.047 |
| LLaMA3.1-8B | 0.238 | 0.210 | 0.028 |
| Qwen2.5-Math-1.5B | 0.356 | 0.227 | 0.129 |
| Claude-3.5 | 0.605 | 0.531 | 0.074 |
| Model | |||
| DS-LLaMA-8B | 0.706 | 0.722 | -0.017 |
| DS-Math-7B | 0.225 | 0.141 | 0.084 |
| QwQ-32B | 0.696 | 0.547 | 0.149 |
| Qwen2.5-7B | 0.362 | 0.130 | 0.232 |
| Marco-o1 | 0.567 | 0.412 | 0.155 |
| DS-Qwen-7B | 0.747 | 0.630 | 0.117 |
| DS-Qwen-1.5B | 0.634 | 0.485 | 0.150 |
| Qwen2.5-3B | 0.402 | 0.209 | 0.193 |
| DS-Qwen-32B | 0.768 | 0.656 | 0.112 |
| DS-Chat | 0.671 | 0.463 | 0.208 |
| Numina-7B | 0.366 | 0.180 | 0.186 |
| DS-Reasoner | 0.810 | 0.750 | 0.060 |
| MetaMath-7B | 0.128 | 0.066 | 0.062 |
| Qwen2.5-Math-7B | 0.355 | 0.191 | 0.164 |
| GPT-4o | 0.564 | 0.497 | 0.067 |
| LLaMA3.1-8B | 0.238 | 0.125 | 0.112 |
| Qwen2.5-Math-1.5B | 0.416 | 0.219 | 0.197 |
| Claude-3.5 | 0.503 | 0.305 | 0.198 |
Appendix F Experiment Setup for Supervised Fine-Tuning
For the fine-tuning experiments, we use Qwen2.5-7B-Instruct and perform full-parameter supervised fine-tuning. All experiments are conducted on two NVIDIA A100 GPUs (80GB memory). The learning rate is set to 1e-5, and training is run for 3 epochs. The training dataset is composed of approximately 100 unreasonable question–answer pairs and 300 well-posed question–answer pairs. The unreasonable samples are constructed from the UMP benchmark, with responses manually edited to clearly identify and explain the flaws through concise, structured reasoning. The well-posed samples are similarly constructed to ensure balanced learning and to mitigate overgeneralization of unreasonable detection behavior.
Appendix G Template for Critical Prompting
Here we show our critical template for solving math problems.
Appendix H Effect of Temperature on Model Behavior
We analyze how different decoding temperatures affect model behavior, focusing on two representative models: Qwen2.5-7B-Instruct (general-purpose) and Qwen-Distill-7B (reasoning-enhanced). Specifically, we evaluate model performance under temperature values of 0.0, 0.2, 0.4, 0.6, 0.8, and 1.0.
As shown in Figure 8, the general-purpose model exhibits relatively stable behavior across temperature settings. In contrast, the reasoning-enhanced model is more sensitive to temperature, particularly when handling unreasonable problems. Notably, at very low temperatures, the reasoning model tends to generate significantly longer outputs, often entering repetitive or overthinking loops. Overall, a temperature of 0.6 yields the most balanced performance across models, aligning with the default setting recommended in the DeepSeek official recommendation.
Appendix I Case Study
I.1 An Example of Overthinking
Table 6 and Table 7 presents representative case of DeepSeek-R1’s behaviors when confronted with unreasonable problems. As illustrated, the model redundantly computes the simple expression “15 2 = 7.5” multiple times throughout its response. Similarly, in the second problem, the model has already identified the misuse of "feet", but is still trying to explain the rationality of the misuse. Table 5 illustrates an example of overthinking behavior in Grok-3-Reasoning. When presented with a simple arithmetic question involving single-digit addition and subtraction, the model engages in excessive reflection, repeatedly generating nearly identical reasoning steps. In this case, the response exceeds 5,600 tokens despite the trivial nature of the task. Moreover, during our experiments, we observed that certain questions could cause the model to hang indefinitely, with some cases taking over 800 seconds without producing a final answer.
For brevity, we omit many highly similar intermediate steps, which do not contribute meaningfully to the final answer. Such repetitive reasoning over a trivial computation is clearly unnecessary and results in substantial computational overhead.
I.2 The unstable of math-specified model
Table 8 illustrates the behavior of Qwen2.5-Math-7B-Instruct when confronted with unreasonable math problems. As shown, the model often becomes paralyzed in the presence of flawed inputs, failing to identify the source of unreasonableness and unable to proceed with a meaningful response.
I.3 Model Unconsciousness of Problem Errors
Tables 9, 10, present examples where the model completely fails to recognize the errors embedded in the questions. These cases suggest that the model entirely overlooks the unreasonableness of the input and proceeds as if the problems were well-posed.
| Unreasonable Problem | Xiao Ming bought a birthday cake for his birthday. He cut the cake into 4 pieces, Xiao Hong ate 3 pieces and Xiao Li ate 2 pieces. How many pieces of cake are left? |
| Unreasonable Part | The total pieces of cake Xiao Ming had initially was 4. If Xiao Hong ate 3 pieces and Xiao Li ate 2 pieces, then they ate 3+2=5 pieces in total. However, since there were only 4 pieces to begin with, it seems like there might be a mistake in the scenario as they couldn’t have eaten 5 pieces from a cake of 4 pieces. So the question is unreasonable. |
| Model Answer | So Xiao Ming bought a birthday cake and cut it into 4 pieces. Then, Xiao Hong ate 3 pieces and Xiao Li ate 2 pieces. I need to find out how many pieces are left.First, I should figure out the total number of pieces initially. It says he cut the cake into 4 pieces, so there are 4 pieces to start with.Next, Xiao Hong ate 3 pieces. So, if there were 4 pieces and Xiao Hong ate 3, I would subtract 3 from 4. That would leave 1 piece.But wait, then Xiao Li ate 2 pieces. If there was only 1 piece left after Xiao Hong ate, how could Xiao Li eat 2 pieces? That doesn’t make sense.Hmm, maybe I need to consider that they ate the pieces at the same time or something. Wait, actually, the problem says Xiao Hong ate 3 pieces and Xiao Li ate 2 pieces, but it doesn’t specify the order or if they ate them simultaneously.Perhaps I should think about the total pieces eaten by both. So, Xiao Hong ate 3 pieces, and Xiao Li ate 2 pieces, so together they ate 3 + 2 = 5 pieces.But originally, there were only 4 pieces. So, if they ate 5 pieces in total, but there were only 4 pieces, that would mean they ate more than what was available, which is impossible.Wait, that can’t be right. [……]Alternatively, perhaps Xiao Hong ate 3 pieces and Xiao Li ate 2 pieces means that together they ate a total of 3 + 2 = 5 pieces, but since there are only 4 pieces, maybe some pieces were eaten by both, but that doesn’t make sense for cake pieces.Wait, unless the pieces are overlapping or something, but cake pieces are typically distinct.This is puzzling. Wait, perhaps it’s about the number of pieces they each consumed, but the pieces are not necessarily whole pieces. For example, maybe they shared some pieces, but the problem says ate 3 pieces, which suggests whole pieces.This is tricky. Maybe I should think about it differently. Perhaps the question is to find out how many pieces are left after both have eaten.But as I calculated earlier, if they ate 3 and 2 respectively, that’s 5 pieces, but there were only 4, so it would be negative, which doesn’t make sense.Alternatively, maybe the problem is that Xiao Ming cut the cake into 4 pieces, but then he had more cakes or something, but that’s not stated.Wait, let’s check the problem again. Xiao Ming bought a birthday cake for his birthday. He cut the cake into 4 pieces, Xiao Hong ate 3 pieces and Xiao Li ate 2 pieces. How many pieces of cake are left?Perhaps it’s a trick question, and the answer is that there are no pieces left, or something like that. But mathematically, if they ate 5 pieces out of 4, it’s impossible.Wait, unless… wait a minute, maybe the pieces that Xiao Li ate include some that Xiao Hong didn’t eat. But since Xiao Hong ate 3 out of 4, there’s only 1 left, and Xiao Li ate 2, but 2 > 1, so still impossible.Alternatively, perhaps ate 3 pieces doesn’t mean 3 whole pieces, but parts of pieces, but the problem says pieces, so likely whole pieces.Wait, here’s another idea: maybe when Xiao Hong ate 3 pieces, those pieces were removed, and then the remaining piece was cut into more pieces, and then Xiao Li ate 2 of those. But the problem doesn’t say that.In fact, the problem is phrased in a way that suggests all actions are on the same set of 4 pieces.Wait, let’s consider the possibility that the cake was cut into 4 pieces, and then Xiao Hong ate 3 pieces, leaving 1 piece, and then that 1 piece was further divided, but again, the problem doesn’t indicate any further cutting.Moreover, even if that were the case, the question is how many pieces of cake are left, which would depend on how many pieces the remaining piece was divided into, but since it’s not specified, that can’t be it.Alternatively, perhaps there’s a cultural or contextual aspect I’m missing, but I doubt it.Wait, [……]This interpretation contradicts the stated consumption. Overlapping Pieces Could Xiao Hong and Xiao Li have eaten some of the same pieces? For example, if Xiao Hong ate pieces A, B, and C, and Xiao Li ate pieces C, D, and one already eaten, the total unique pieces consumed might be less than 5. But with only 4 pieces (say, A, B, C, D), and Xiao Hong eating 3 (A, B, C), Xiao Li’s 2 pieces (e.g., C, D) overlap on C, suggesting a total of 4 unique pieces eaten. Yet, cake pieces are typically distinct—once eaten, a piece isn’t available for someone else unless shared fractionally, which isn’t implied here. Misstated Numbers Maybe there’s a typo. If Xiao Hong ate 1 piece and Xiao Li ate 2 pieces, the total eaten would be 3, leaving 4 - 3 = 1 piece. [……]If we assume they consumed all 4 pieces between them, the maximum possible pieces eaten is 4, leaving 0. But Xiao Hong’s 3 plus Xiao Li’s 2 is 5, exceeding 4, suggesting the problem’s numbers may not align with a simple subtraction model. Step 6: Hypothesize a Logical OutcomeIn many word problems, especially those involving tangible items like cake, the intent is for the numbers to work sensibly. The phrase “how many pieces are left” implies some remainder is possible, but here, eating 5 out of 4 means at least all pieces are gone. Let’s test the boundary:- Maximum pieces available: 4- Minimum pieces eaten (if constrained): 4 (all pieces consumed)- Pieces left: 4 - 4 = 0 This assumes they somehow ate only 4 pieces despite the stated 5, perhaps implying Xiao Li ate 1 piece after Xiao Hong’s 3, but the problem insists on 2 pieces for Xiao Li. The most consistent resolution, given no additional cake, is that all pieces were eaten, and the extra “1 piece” is impossible, leaving 0. Step 7: Conclusion After exploring sequential eating, overlaps, and errors, the simplest resolution is that the cake’s 4 pieces were fully consumed, and the stated 5 pieces eaten exceeds the total, leaving none. This might indicate a flaw in the problem’s design (e.g., a typo), but working strictly with the given text—“4 pieces total, 3 eaten by Xiao Hong, 2 by Xiao Li”—the logical endpoint, avoiding negative pieces, is 0 pieces left. They couldn’t eat more than 4, so all pieces are gone.**Final Answer: 0** |
| Column 1 | Column 2 |
| Unreasonable Problem | Jenna and her mother picked some apples from their apple farm. Jenna picked half as many apples as her mom. If her mom got 15 apples, how many apples did they both pick? |
| Unreasonable Part | This question is unreasonable because picking half an apple is not practical in real-life scenarios (as apples are typically picked whole). In this case, if we adhere strictly to the question, they picked a total of 15 (mom’s apples) + 7.5 (Jenna’s apples) = 22.5 apples. The only unreasonable aspect might be the fractional apple, as it suggests that an apple was divided, which is not typical when picking apples from a tree. |
| Model Answer | Okay, so Jenna and her mother picked apples from their apple farm. The problem says Jenna picked half as many apples as her mom. And her mom got 15 apples. The question is, how many apples did they both pick together? Let me try to break this down step by step. First, I need to figure out how many apples Jenna picked. Since Jenna picked half as many as her mom, and her mom picked 15 apples, I can represent Jenna’s apples as half of 15. So, mathematically, that would be 15 divided by 2. Let me write that out: 15 ÷ 2. Hmm, doing that division, 15 divided by 2 is 7.5. Wait a second, can you have half an apple? That seems a bit odd because you can’t really pick half an apple in real life. Maybe the problem expects a whole number? Let me check the problem again. It says Jenna picked half as many apples as her mom. Her mom got 15 apples. Hmm, maybe the answer is 7.5 apples, but since apples are whole, maybe we need to round? But the problem might just be using numbers that divide evenly. Wait, 15 divided by 2 is 7.5. But maybe I made a mistake here. Let me verify. If her mom picked 15 apples, and Jenna picked half as many, then yes, it’s 15 divided by 2. But 15 is an odd number, so dividing by 2 gives a decimal. But apples are countable items; you can’t have half. So perhaps there’s a mistake in the problem? Or maybe I need to assume that half can be a fraction here even though it’s not practical. Let me think. The problem might just be expecting a fractional answer even if it’s not a whole number, for the sake of the math problem. Maybe in the context of this problem, fractions are acceptable even if they aren’t in real life. Alternatively, maybe the numbers are supposed to work out evenly, and I need to check if I misread the problem. Let me reread it: "Jenna picked half as many apples as her mom. If her mom got 15 apples, how many apples did they both pick?" So, mom has 15, Jenna has half, which is 7.5. Together, that would be 15 + 7.5 = 22.5 apples. But again, in reality, you can’t have half an apple. Maybe the question allows decimal answers, or perhaps there’s an error in the problem. […] Jenna picked half as many apples as her mom." So Jenna’s apples = (1/2) * mom’s apples. Mom’s apples are 15, so Jenna’s apples are (1/2)*15 = 7.5. Then total apples would be 15 + 7.5 = 22.5. But that’s 22 and a half apples. Since we can’t have half apples, perhaps the problem is intended to have whole numbers, and there’s a typo. […] Alternatively, maybe there’s a mistake in the problem’s numbers. If the mom had 14 apples, then Jenna would have 7, total 21. If the mom had 16, Jenna 8, total 24. But with 15, it’s 7.5. So perhaps the problem is testing the understanding of fractions even in context where it might not make practical sense. So maybe the answer is 22.5. But since you can’t have half an apple, maybe the answer is 23, rounding up? But that’s not specified. Alternatively, maybe the answer is 22, rounding down? But again, the problem doesn’t say to round. Alternatively, maybe the problem is expecting a mixed number, like 22 1/2. But in terms of the answer, they might want it in a box as or . But in most math problems like this, unless specified, they might expect a fractional answer. […] The problem says mom has 15. So that interpretation is not correct. Because if it’s a ratio, then total would be different. But the problem explicitly states the mom has 15. So that’s different. Alternatively, maybe the problem is saying "Jenna picked half as many apples as her mom [did]", which is the same as Jenna’s apples = 1/2 * mom’s apples. So that’s 7.5. Therefore, total is 22.5. […] Jenna picked half as many apples as her mother. Her mother picked 15 apples. To find the total number of apples they both picked, we first determine how many apples Jenna picked. Jenna’s apples are calculated as: Next, we add the number of apples picked by Jenna and her mother to find the total: Thus, the total number of apples Jenna and her mother picked is . |
| Column 1 | Column 2 |
| Unreasonable Problem | Compute in your head. |
| Unreasonable Part | Set the units of to feet and the other terms to pure numbers. Inconsistent units prevent meaningful addition operations. |
| Model Answer | Okay, so I need to compute 99 squared feet plus 99 plus 1. Hmm, let me think. First, the problem is asking for 99² feet + 99 + 1. Wait, is the feet unit important here? The question mentions compute 99² feet + 99 + 1, but all the terms except maybe the first one don’t have units specified. Maybe it’s a typo or maybe they are all in feet? Hmm, but usually, you can’t add feet to a unitless number. Hmm, maybe the units are just there for confusion? Let me check again.Wait, the original problem is written as: Compute feet + 99 + 1 So the first term is 99 squared feet, and the other terms are just numbers without units. But you can’t add feet to unitless numbers. That doesn’t make sense dimensionally. So maybe the problem is actually written with a typo? Or perhaps the + 99 + 1 are also in feet? Maybe the question meant 99² square feet plus 99 square feet plus 1 square foot? Because otherwise, adding feet to pure numbers is impossible.Alternatively, maybe the entire expression is unitless, and the feet was a mistake. The problem says compute feet + 99 +1 in your head. Maybe it’s supposed to be 99 squared plus 99 plus 1? So 99² + 99 + 1. Then the answer would be a number, not in feet. That seems more plausible. Maybe the mention of feet is a red herring or a typo. Let me check if that’s possible.Alternatively, maybe it’s a play on words. Like 99 square feet plus 99 feet plus 1 foot.[……] So maybe the problem is intended to be 99 squared plus 99 plus 1, with feet being a distractor. Let me assume that. Because otherwise, the problem doesn’t make mathematical sense.So assuming the problem is 99² + 99 + 1, compute that in your head. Let me compute that. So 99 squared is… Hmm, 99 times 99. How can I compute that quickly? Well, 99 is 100 minus 1, so 99 squared is . Which is 100² - 2*100*1 + 1² = 10,000 - 200 + 1 = 9,801. Right? Let me check: 100 squared is 10,000, subtract 200 gives 9,800, plus 1 is 9,801. Yep. So 99² is 9,801. Then adding 99 gives 9,801 + 99. Hmm, 9,801 + 99 is 9,900. Then adding 1 gives 9,901. So total is 9,901. So the answer would be 9,901. But wait, if the original problem had units, like 99² square feet plus 99 feet plus 1 foot, then the answer would be nonsensical. But since the problem is written as , maybe they just want the numerical value, disregarding units?[……] Hmm. Maybe I need to clarify.Alternatively, the problem might be using as a notation for 99 feet squared, which is ambiguous. Because feet squared can mean square feet or (feet) squared. But in any case, given the problem as stated, I think the most plausible interpretation is that it’s a mathematical expression where all terms are unitless, except the first one has feet attached. But since you can’t add them, maybe the units are a mistake. So the problem is likely 99 squared plus 99 plus 1, which equals 9,901. So I think that’s the intended answer. Therefore, the answer is . Alternatively, if units are considered, and all are in feet, the answer is 9,901 feet, so feet. But since the problem says compute… in your head, probably expecting a numerical answer, so 9,901. Yeah, I think that’s it.**Final Answer** |
| Column 1 | Column 2 |
| Unreasonable Problem | Two candidates are running for class representative at Sarai’s school. If the winner got 5/6 of the votes and the total number of students who voted in the class was 80, calculate the number of votes the loser got. |
| Unreasonable Part | This question is unreasonable because it implies that the number of votes received by both candidates does not account for the total number of votes cast. If the winner got 5/6 of the votes from 80 students, this means the winner received approximately 66.67 votes (which cannot be a fraction as a person cannot cast a fraction of a vote). The problem then asks to calculate the number of votes the loser got, but since there are only 80 votes in total and 66.67 have been allocated to the winner, this leaves 13.33 votes for the loser. However, since votes are discrete and cannot be fractional, having a fraction of a vote is not possible, making the question unreasonable as it cannot accurately reflect a real-life voting scenario where each voter casts exactly one whole vote. |
| Model Answer | To determine the number of votes the loser received, we need to follow these steps: 1. Calculate the number of votes the winner received. 2. Subtract the number of votes the winner received from the total number of votes to find the number of votes the loser received. First, let’s calculate the number of votes the winner received. The winner got of the total votes, and the total number of votes is 80. Therefore, we multiply star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star[……]star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star star. |
| Column 1 | Column 2 |
| Unreasonable Problem | A regular hexagon can be divided into six equilateral triangles. If the perimeter of one of the triangles is 21 square inches, what is the perimeter, in inches, of the regular hexagon? |
| Unreasonable Part | The circumference should be expressed in length units (such as inches), but the incorrect use of area units (square inches) in the title leads to logical contradictions in units and cannot be calculated numerically. |
| Model Answer | To determine the perimeter of the regular hexagon, we start by analyzing the given information about the equilateral triangle. We know that the perimeter of one of the equilateral triangles is 21 inches. Since an equilateral triangle has three equal sides, we can find the length of one side of the triangle by dividing the perimeter by 3: A regular hexagon can be divided into six equilateral triangles, and each side of the hexagon is equal to the side length of these triangles. Therefore, the side length of the hexagon is also 7 inches. The perimeter of a regular hexagon is the sum of the lengths of its six sides. Since all sides are equal, the perimeter is: Thus, the perimeter of the regular hexagon is (). |
| Column 1 | Column 2 |
| Unreasonable Problem | What is the least positive integer multiple of 30 that can be written with only the digit 2? |
| Unreasonable Part | Multiples of 30 must end in 0, and the problem restricts the use of only the number 2, resulting in no eligible numbers (unable to satisfy the requirement of ending in 0). |
| Model Answer | To find the least positive integer multiple of 30 that can be written with only the digit 2, we need to consider the properties of multiples of 30. A number is a multiple of 30 if and only if it is a multiple of both 3 and 10. This means the number must end in 0 and the sum of its digits must be a multiple of 3. Since the number can only contain the digit 2, it must end in 0. However, a number consisting only of the digit 2 cannot end in 0. Therefore, we need to consider the possibility of a number that is a multiple of 30 and can be written with only the digit 2, but not necessarily ending in 0. This means we need to find a number that is a multiple of 3 and can be written with only the digit 2, and then check if it is a multiple of 10. Let’s start by checking the smallest numbers consisting only of the digit 2: - 2 (not a multiple of 30) - 22 (not a multiple of 30) - 222 (sum of digits = 2 + 2 + 2 = 6, which is a multiple of 3, but does not end in 0) - 2222 (sum of digits = 2 + 2 + 2 + 2 = 8, which is not a multiple of 3) - 22222 (sum of digits = 2 + 2 + 2 + 2 + 2 = 10, which is not a multiple of 3) - 222222 (sum of digits = 2 + 2 + 2 + 2 + 2 + 2 = 12, which is a multiple of 3, and ends in 2, so it is not a multiple of 10) We need to find a number that is a multiple of 3 and ends in 0. The smallest such number is 2222220, but we need to check if there is a smaller number. Since 2222220 is the smallest number that meets the criteria (it is a multiple of 3 and ends in 0), we can conclude that it is the least positive integer multiple of 30 that can be written with only the digit 2. Thus, the answer is (). |
| Column 1 | Column 2 |
| Unreasonable Problem | John has 3 boxes. Each box is 5 grams by 6 grams by 4 grams. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? |
| Unreasonable Part | Describing the size (length, width, and height) of a box using the unit of mass "grams" mixed with the unit of volume "inches" leads to a logical contradiction in the unit. |
| Model Answer | First, I need to determine the inner dimensions of each box by subtracting twice the wall thickness from each external dimension. The external dimensions are 5 grams by 6 grams by 4 grams, and the walls are 1 inch thick. However, there’s a unit inconsistency since the dimensions are in grams and the wall thickness is in inches. I should convert the wall thickness to grams to maintain consistent units. Assuming 1 inch is approximately 25.4 grams, the inner dimensions would be:- Length: 5 grams - 2 * 25.4 grams = -45.8 grams- Width: 6 grams - 2 * 25.4 grams = -44.8 grams- Height: 4 grams - 2 * 25.4 grams = -46.8 grams Since the inner dimensions cannot be negative, this indicates that the wall thickness is too large relative to the external dimensions, making the inner volume zero. Therefore, the total inner volume of all three boxes is zero.</think>To determine the total inner volume of all 3 boxes, we’ll follow these steps:1. Understand the Dimensions: External Dimensions of Each Box: 5 grams × 6 grams × 4 grams Wall Thickness: 1 inch 2. Convert Units for Consistency: Note: The external dimensions are given in grams, while the wall thickness is in inches. To proceed, we need consistent units. Assuming 1 inch = 25.4 grams. Wall Thickness in Grams: 1 inch = 25.4 grams 3. Calculate Inner Dimensions: Length: 5 grams - 2 × 25.4 grams = 5 - 50.8 = -45.8 grams Width: 6 grams - 2 × 25.4 grams = 6 - 50.8 = -44.8 grams Height: 4 grams - 2 × 25.4 grams = 4 - 50.8 = -46.8 grams 4. Analyze the Results: - The inner dimensions are negative, which is not physically possible. This indicates that the wall thickness is too large relative to the external dimensions, resulting in no inner volume.5. Conclusion: - Since the inner dimensions cannot be negative, the inner volume of each box is 0. - Therefore, the total inner volume of all 3 boxes is also . |
| Column 1 | Column 2 |
| Unreasonable Problem | Cars have lined up on the motorway. several cars drive through in the first 15 minutes of the traffic jam, then 20 more cars drive through in the remaining 15 minutes of the jam. 15 cars from the line take an exit so they don’t have to drive through the traffic jam. If there were originally 30 cars on the motorway, how many cars drove through the traffic jam in the first 15 minutes? |
| Unreasonable Part | This question is unreasonable. The total number of cars that drove through and took an exit (several + 20 + 15) is more than the original number of cars on the motorway (30) |
| Model Answer | Let me help solve this step by step.1) First, let’s identify the key information:Initially there were 30 cars on the motorway15 cars took an exit to avoid the jam20 cars drove through in the last 15 minutesWe need to find how many cars drove through in the first 15 minutes2) Let’s solve:Out of the initial 30 cars: - 15 cars took the exit - This means 30 - 15 = 15 cars remained to go through the jam3) We know:20 cars went through in the last 15 minutes15 cars from the original group had to go throughTotal cars that went through = First 15 min + Last 15 min (20)4) Therefore:If 15 total original cars went throughAnd 20 went through in the last 15 minutesThen in the first 15 minutes: 15 - 20 = -5 cars from the original group went through5) This means that all cars that went through in the first 15 minutes must have been new cars (not from the original 30)The answer is cars drove through the traffic jam in the first minutes. |
| Column 1 | Column 2 |
| Unreasonable Problem | A salesman bought a case of 48 sneakers for $576. He sold 17 of them for $20 at a flash sale on Monday, and the remaining 35 were sold to a department store for $25 each. How much money, in dollars, was the salesman’s profit? |
| Unreasonable Part | The total sold quantity (17+35=52 pairs) exceeds the total purchased quantity (48 pairs), resulting in a contradiction between the quantity before and after, and the remaining quantity cannot be established. |
| Model Answer | To determine the salesman’s profit, we’ll follow these steps:1. Calculate the Total Cost: The salesman bought 48 sneakers for $576. 2. Calculate the Revenue from the Flash Sale: He sold 17 sneakers at $20 each. 3. Calculate the Revenue from the Department Store: The remaining 35 sneakers were sold at $25 each. 4. Calculate the Total Revenue: Add the revenues from both sales. 5. Calculate the Profit: Subtract the total cost from the total revenue. Final Answer: |
| Column 1 | Column 2 |
| Unreasonable Problem | Four books are arranged on a shelf. The first book is 3100 mm thick while the second book is 50 mm thick. The third book is 5 mm less thick than the second book, and the fourth book is twice as thick as the first book. What is the total thickness of the four books? |
| Unreasonable Part | The thickness of the first book is 3100 millimeters (3.1 meters), far exceeding the reasonable thickness of real books, resulting in an unrealistic calculation of the total thickness. |
| Model Answer | <think>First, I identify the thickness of each book individually.The first book is 3100 mm thick.The second book is 50 mm thick.The third book is 5 mm less than the second book, so that’s 50 mm minus 5 mm, which equals 45 mm.The fourth book is twice as thick as the first book, so that’s 2 multiplied by 3100 mm, which equals 6200 mm.Next, I add up the thicknesses of all four books to find the total thickness.Adding 3100 mm, 50 mm, 45 mm, and 6200 mm together gives a total thickness of 9395 mm.</think>Let’s determine the total thickness of the four books step by step.1. First Book: - Thickness = 3100 mm2. Second Book: - Thickness = 50 mm3. Third Book: - It is 5 mm less thick than the second book. - Thickness = 50 mm - 5 mm = 45 mm4. Fourth Book: - It is twice as thick as the first book. - Thickness = 2 × 3100 mm = 6200 mm5. Total Thickness: - Add the thicknesses of all four books together. - Total Thickness = 3100 mm + 50 mm + 45 mm + 6200 mm = mm" |
| Column 1 | Column 2 |
| Unreasonable Problem | Uncle Ben has four horses. Each horse can consume five pounds of oats per meal. He feeds his horses twice a day. If one bag contains 50 pounds of oats, how many bags of oats does he need to feed his five horses for five days? |
| Unreasonable Part | At the beginning of the question, it is clearly stated that there are 4 horses, but the problem requires calculating the amount of oats needed to feed 5 horses, which leads to contradictory conditions and makes it impossible to determine whether 4 or 5 horses should be used for calculation. |
| Model Answer | First, determine the total amount of oats needed per horse per day. Each horse consumes 5 pounds of oats twice a day, so that’s 10 pounds per horse per day.Next, calculate the total oats needed for all five horses per day by multiplying the oats per horse by the number of horses: 10 pounds/horse/day * 5 horses = 50 pounds per day.Then, find out the total oats needed for five days by multiplying the daily requirement by five: 50 pounds/day * 5 days = 250 pounds.Finally, determine how many bags of oats are needed by dividing the total oats required by the weight of each bag: 250 pounds / 50 pounds per bag = 5 bags.</think>Let’s break down the problem step by step to find out how many bags of oats Uncle Ben needs.1. C alculate the daily oats consumption for one horse: - Oats per meal per horse: 5 pounds - Number of meals per day: 2 - Total oats per day per horse: 2. Calculate the total oats needed for all five horses per day: - Number of horses: 5 - Total oats per day for all horses: 3. Calculate the total oats needed for five days: - Number of days: 5 - Total oats for five days: 4. Determine the number of bags needed: - Weight of one bag: 50 pounds - Number of bags needed: Final Answer: |
Appendix J Example output by different LLM
In this section, we will demonstrate how various models perform on unreasonable problems using the conventional Chain of Thought approach compared to adding a critical instruction method. Tables 16 display the outputs from Qwen2.5-7B-Instruct using two different prompts. It is evident that the outputs generated by adding critical instruction exhibit marked enhancements.
| Unreasonable problem | Baseline output | Critical output |
| Timothy’s locker is 24 inches. Zack’s locker is half as big as Timothy’s locker. Peter’s locker is 1/4 as big as Zack’s locker. How big is Peter’s locker in cubic inches? | To determine the size of Peter’s locker, we need to follow the given relationships step by step.1. **Determine the size of Zack’s locker:** - Timothy’s locker is 24 inches. - Zack’s locker is half as big as Timothy’s locker. 2. **Determine the size of Peter’s locker:** - Peter’s locker is as big as Zack’s locker. Since the problem asks for the size of Peter’s locker in cubic inches, we need to assume that the lockers are cubes (as the problem does not specify otherwise).3. **Calculate the volume of Peter’s locker:** - The side length of Peter’s locker is 3 inches. - The volume of a cube is given by the formula , where is the side length. Thus, the size of Peter’s locker in cubic inches is . | Let’s analyze the problem step by step to determine the size of Peter’s locker.1. **Timothy’s Locker:** - Timothy’s locker is given as 24 inches. However, the problem does not specify the dimensions (length, width, height) of the locker. For a locker to be 24 inches, it would need to be a one-dimensional measurement, which is not typical for a locker. Typically, a locker would have three dimensions (length, width, and height).2. **Zack’s Locker:** - The problem states that Zack’s locker is half as big as Timothy’s locker. If we assume half as big refers to volume, we need to know the dimensions of Timothy’s locker to proceed. Without this information, we cannot accurately determine the volume of Zack’s locker.3. **Peter’s Locker:** - The problem further states that Peter’s locker is 1/4 as big as Zack’s locker. Again, without knowing the dimensions of Zack’s locker, we cannot determine the volume of Peter’s locker.Given the lack of specific dimensions for each locker, the problem is unreasonable. We cannot determine the size of Peter’s locker in cubic inches without additional information about the dimensions of the lockers.**Conclusion:**The problem is unreasonable because it does not provide sufficient information to determine the dimensions of the lockers, which are necessary to calculate the volume. Therefore, we cannot provide a specific answer for the size of Peter’s locker in cubic inches.Insufficient information |
| James decides to buy birthday candles for his 2 sons. One of them is 12 and the other is 15 years younger. A pack of 5 candles costs $3. How much does James spend on candles? | To determine how much James spends on candles, we need to follow these steps:1. **Determine the ages of James’s sons:** - The first son is 12 years old. - The second son is 15 years younger than the first son. - Therefore, the age of the second son is: - Since age cannot be negative, we need to re-evaluate the problem. The second son being 15 years younger than the first son implies he is 12 - 15 = -3, which is not possible. This suggests a misinterpretation. Let’s assume the second son is 15 years younger in absolute terms, meaning he is: (which is not possible, so we consider the absolute value) - Correct interpretation: The second son is 12 - 15 = -3, which is not possible. Let’s assume the second son is 15 years younger in absolute terms, meaning he is: (which is not possible, so we consider the absolute value)3. **Determine the number of packs needed:** - Each pack contains 5 candles. - The total number of candles needed is 12 + 12 = 24. - The number of packs needed is: - Since James cannot buy a fraction of a pack, he needs to buy 5 packs.4. **Calculate the total cost:** - Each pack costs $3. - The total cost is: Therefore, the total amount James spends on candles is . | Let’s break down the problem step by step to ensure it is reasonable and correctly solved.1. **Determine the ages of the sons:** - The first son is 12 years old. - The second son is 15 years younger than the first son. - Therefore, the second son’s age is .The age of -3 years is not reasonable for a human being. This indicates that there is an error in the problem statement. It is not possible for one son to be 15 years younger than another son who is 12 years old.2. **Revisiting the problem:** - If we assume the problem meant the second son is 15 years younger in a more reasonable context, we should consider the possibility of a typo or misinterpretation. For example, if the second son is 15 years younger in a different context (like a different age difference), we would need to clarify the problem. Since the problem as stated is unreasonable, we cannot proceed with the calculation as intended. The issue here is the logical inconsistency in the ages provided.**Conclusion:**The problem is unreasonable due to the negative age of the second son. Therefore, we cannot determine how much James spends on candles based on the given information.Unreasonable problem |
| A plane travels 1200 miles in 3 miles. At the same rate, how many additional hours would it take to travel an additional 2000 miles? | To solve this problem, we need to determine the speed of the plane and then use that speed to find out how long it would take to travel an additional 2000 miles.Step 1: Determine the speed of the plane.The plane travels 1200 miles in 3 hours. Therefore, the speed of the plane is:Step 2: Calculate the time it would take to travel an additional 2000 miles.Using the speed we found, we can calculate the time it would take to travel 2000 miles:Therefore, it would take an additional hours to travel 2000 miles. | The problem as stated contains an error. The phrase 3 miles is likely meant to be 3 hours, but even then, the initial statement A plane travels 1200 miles in 3 miles is not logically consistent. Let’s break it down step by step:1. **Initial Statement Analysis:** - The statement A plane travels 1200 miles in 3 miles is incorrect. It should be A plane travels 1200 miles in 3 hours.2. **Corrected Statement:** - If the plane travels 1200 miles in 3 hours, we can calculate the speed of the plane. - Speed = Distance / Time - Speed = 1200 miles / 3 hours = 400 miles per hour.3. **Additional Travel Calculation:** - We need to find out how many additional hours it would take to travel an additional 2000 miles at the same speed. - Time = Distance / Speed - Time = 2000 miles / 400 miles per hour = 5 hours.Therefore, the plane would take an additional hours to travel 2000 miles at the same rate. |
Appendix K Evaluation Prompt Details
We use an LLM-as-a-judge setup to assess whether a model response reasonably identifies the unreasonableness in the modified question. The full evaluation prompt is shown below. For each dataset, we design corresponding in-context examples to better assess the model’s performance.
Appendix L Prompt for Unreasonable Problem Generation
We present the detailed prompt used for generating unreasonable math problems via LLMs. Note that the actual prompt is written in Chinese, and the English version shown below is a direct translation for presentation purposes.
Appendix M Examples in UMP benchmark
In this section we will show one unreasonable problem for each category in our benchmark in table 17. “Answer” refers to the solution to the original problem. “New Question” denotes the artificially generated question that is designed to be unreasonable. Among them, the part that becomes unreasonable by modification is already represented in bold. Accompanying “Explanation” provides rationale for the question’s lack of reason, serving as a guide for evaluators to scrutinize the original problem’s construction. Our intention is that this explanation aids test creators in reviewing and refining their problems to ensure logical consistency.
Appendix N Query format for Pre-validation experiment
Our prompt is as follows: Please decide whether the following statement is reasonable:{d}.Where “d” is the description shown in the following table
| Original question | Answer | Type | New question | Explanation |
| In a race with 240 Asians, 80 were Japanese, and the rest were Chinese. If the number of boys on the Chinese team was 60, how many girls were on the Chinese team? | 100 | Incorrect Assumptions | In a race with 240 Asians, 80 were Japanese, 120 were South Korean, and the rest were Chinese. If the number of boys on the Chinese team was 60, how many girls were on the Chinese team? | This question is unreasonable because the total number of Chinese racers is implied to be less than the number of boys on the Chinese team, which is a mathematical impossibility. Specifically, the total number of racers is 240, with 80 Japanese and 120 South Korean, leaving 40 Chinese racers (240 - 80 - 120 = 40). However, the question then states that there are 60 boys on the Chinese team, which cannot be true as this would require the total number of Chinese racers to be at least 60, contradicting the earlier calculation that there are only 40 Chinese racers in total. Therefore, the question is based on an incorrect premise that makes it logically and mathematically flawed. |
| A car is on a road trip and drives 60 mph for 2 hours, and then 30 mph for 1 hour. What is the car’s average speed in mph during this trip? | 50 | Misinterpretation of Units | A car is on a road trip and drives 60 square miles per hour for 2 hours, and then 30 cubic miles per hour for 1 hour. What is the car’s average speed in mph during this trip? | This question is unreasonable because it misuses units of measurement by introducing "square miles per hour" and "cubic miles per hour," which are nonsensical. Speed should be measured in units of distance over time (e.g., miles per hour), not area or volume over time. |
| Brittany and her mom go to the museum. The cost of admission is $12 for adults and $10 for children. Brittany’s mom gives the cashier money for 1 child ticket and 1 adult ticket. If she received $8 in change, how much money, in dollars, did she give the cashier? | 30 | Inconsistent Conditions | Brittany and her mom go to the museum. The cost of admission is $12 for adults and $10 for children. Brittany’s mom gives the cashier money for 1 child ticket and 1 adult ticket. If she received $8 in change, how much money, in dollars, did she give the cashier if she paid for 3 adult tickets and 2 children tickets? | This question is unreasonable because it contradicts itself. The first part of the question states that Brittany’s mom paid for only 1 adult ticket and 1 child ticket. However, the latter part of the question contradicts this by assuming that she paid for 3 adult tickets and 2 children tickets. This makes the question impossible to answer accurately, as we are given conflicting information. |
| Carl buys ten packs of cookies. Each pack of cookies has six cookies inside. Each cookie cost $0.10. How much change does Carl receive if he pay with a $10 bill? | 4 | Illogical Scenarios | Carl buys ten packs of cookies. Each pack of cookies has six cookies inside. Each cookie costs $0.10. How many kittens does Carl have if he pays with a $10 bill? | This question is unreasonable because it presents an illogical situation. The number of kittens Carl has is completely unrelated to the amount of money he spent on cookies or the payment method he used. It is not possible to deduce the number of kittens Carl has based on the information given about his cookie purchase. Thus, this question cannot be logically or reasonably answered with the provided information. |
| Misha picks out 4 blouses from the 30% off rack. The regular price for each blouse is $20. How much is the total cost of the discounted blouses? | 56 | Undefined Variables | If Misha picks out some blouses from the 30% off rack, how much is the total cost of the discounted blouses? | This question is unreasonable because it lacks the specific number of blouses Misha picked, which is crucial to calculate the total cost of the discounted blouses. Without knowing the quantity of blouses chosen, it’s impossible to determine the total cost. |

### v6

@@
-Large Language Models Are Struggle to Cope with Unreasonability in Math Problems
+Large Language Models Struggle with Unreasonability in Math Problems
 Abstract
-Recent research have demonstrated LLMs’ impressive performance in math and reasoning.
-However, the capacity of LLMs to address math problems under unconventional conditions, such as internal inconsistencies and flawed assumptions, remains largely unexplored.
-In this paper, we propose a novel benchmark Unreasonable Math Problem (UMP) designed to assess LLMs’ ability to recognize and respond to unreasonability in math problem.
-The benchmark consists of a carefully curated collection of unreasonable math questions across diverse types.
-Based on extensive experiments covering 19 LLMs, we observe that even state-of-the-art models such as GPT-4o achieve only limited performance of 0.6 in UMP, while reasoning models such as DeepSeek-R1 are prone to overthinking and unstable.
-We further explore strategies for improving the recognition of unreasonable inputs, shedding light on both the possibility and limitations of LLMs in this challenging setting.
-Large Language Models Are Struggle to Cope with Unreasonability
-in Math Problems
+Large Language Models (LLMs) have shown remarkable success on a wide range of math and reasoning benchmarks.
+However, we observe that they often struggle when faced with unreasonable math problems.
+Instead of recognizing these issues, models frequently proceed as if the problem is well-posed, producing incorrect answers or falling into overthinking and verbose self-correction.
+To systematically investigate this overlooked vulnerability, we propose the Unreasonable Math Problems (UMP) benchmark, designed to evaluate LLMs’ ability to detect and respond to unreasonable math problem statements.
+Based on extensive experiments covering 19 LLMs, we find that even state-of-the-art general models like GPT-4o achieve only a score of 0.6 on UMP.
+While reasoning models such as DeepSeek-R1 demonstrate a higher sensitivity to unreasonable inputs, this often comes at the cost of generating overly long and meaningless responses that fail to converge.
+We further explore prompting and fine-tuning methods, which offer partial improvements but also introduce trade-offs, shedding light on both the potential and limitations of LLMs in this challenging setting.
+Large Language Models Struggle with Unreasonability in Math Problems
 Jingyuan Ma1 , Damai Dai1 , Zihang Yuan2, Rui Li1, Weilin Luo3, Bin Wang3, Qun Liu3, Lei Sha2, Zhifang Sui1 1National Key Laboratory for Multimedia Information Processing, School of Computer Science, Peking University 2Institute of Artificial Intelligence, Beihang University 3Huawei Noah’s Ark Lab, China {mjy@stu.pku.edu.cn}
 1 Introduction
-Large language models (LLMs) have recently achieved remarkable progress in math and reasoning, particularly with the emergence of reasoning-enhanced models like O1 and R1 DeepSeek-AI (2025), demonstrate increasingly structured and deliberate reasoning capabilities.
-As a result, current LLMs have surpassed earlier baselines and achieved over 90% accuracy on math benchmarks like GSM8K Cobbe et al.
-(2021a) and MATH Hendrycks et al.
-(2021).
-However, despite this impressive performance on challenging math problems, we find that these models often struggle to recognizing when a math question is unreasonable.
-Even when the irrationality, such as a logical contradiction or a violation of basic constraints, is immediately obvious to humans.
-LLMs frequently follow the misleading question setup, producing confident yet nonsensical answers, or becoming trapped in cycles of excessive reasoning that fail to converge.
-This counter-intuitive behavior raises serious concerns about the reliability in their practical applications such as automated tutoring Kasneci et al.
+Large language models (LLMs) have recently shown impressive performance on advanced mathematical reasoning tasks, especially on benchmarks like MATH (Hendrycks et al., 2021) and AIME24MAA (2024).
+However, we find that these models often fail to detect logical flaws or unreasonable assumptions in math problems, treating them as if they were well-posed.
+Instead of flagging such issues, they tend to generate confident yet nonsensical answers, or fall into endless reasoning loops without reaching a valid conclusion.
+This counter-intuitive behavior raises serious concerns about their reliability in real-world applications such as automated tutoring Kasneci et al.
 (2023), early education Zhang et al.
-(2024b), or open-domain problem solving Lin and Chen (2023).
+(2024b), and open-domain problem solving Lin and Chen (2023), where misleading answers to unreasonable questions can undermine trust and lead to negative outcomes.
 To enable a comprehensive analysis of how LLMs behave when confronted with mathematically unreasonable inputs, we introduce the Unreasonable Math Problems (UMP) benchmark.
-We take questions from GSM8K Cobbe et al.
-(2021a), MATH Hendrycks et al.
-(2021), AIME24 MAA (2024), and AMC23 to generated unreasonable variants.
-Specifically, we utilize GPT-4o to apply minimal edits that introduce specific forms of unreasonability while preserving the original questions structure.
-These edits are guided by a set of rule-based transformation templates.
-All generated questions are verified to ensure that they are genuinely unreasonable in a mathematical or logical sense, and that they remain similar to the original questions in length, topic, and surface form.
-This careful construction allows us to attribute differences in model behavior specifically to the presence of unreasonableness, rather than to confounding changes in difficulty.
-Based on UMP benchmark, we conducted extensive experiments on various models spanning three categories: general-purpose models (e.g., GPT-4oOpe (2023)), reasoning-enhanced models (e.g., DeepSeek-R1DeepSeek-AI (2025)), and math-specialized models (e.g., Qwen-MathYang et al.
+We construct UMP by minimally editing questions from existing math datasets, including MATH Hendrycks et al.
+(2021), AIME24 MAA (2024), AMC23, and GSM8K Cobbe et al.
+(2021a) to create unreasonable variants that contain logical inconsistencies, missing assumptions, or ill-defined objectives.
+These edits are guided by rule-based transformation templates and executed using DeepSeek-R1, while preserving the original problem’s structure, topic, and surface form.
+All generated questions are manually verified to ensure they are mathematically flawed yet still natural.
+To evaluate model behavior, we present models with both the original and its corresponding unreasonable version.
+This paired setup enables us to directly attribute behavioral differences to the presence of unreasonableness, isolating it from other factors such as question length or difficulty.
+Based on the UMP benchmark, we conduct extensive experiments on various models spanning three categories: general-purpose models (e.g., GPT-4oOpenAI (2024)), reasoning models (e.g., DeepSeek-R1DeepSeek-AI (2025)), and math-specialized models (e.g., Qwen-MathYang et al.
 (2024a)).
 In addition, we analyze LLMs’ failure patterns according to tokens repetition, reflection frequency, and token entropy.
-We find: 1) general-purpose models often proceed confidently without recognizing the unreasonableness of the question; 2) reasoning-enhanced models tend to overthink and fall into excessive self-correction; 3) math-specialized models may fail to initiate reasoning when confronted with unreasonable premises.
-We further explored several mitigation strategies but found that none could robustly resolve these failure modes without introducing trade-offs, such as decreased performance on standard inputs.
+We find: 1) general-purpose models often proceed confidently without recognizing the unreasonableness of the question; 2) reasoning models tend to overthink and fall into excessive self-correction; 3) math-specialized models may fail to initiate reasoning when confronted with unreasonable premises.
+We further explore several mitigation strategies but find that none could robustly resolve these failure modes without introducing trade-offs, such as decreased performance on standard inputs, highlighting the need for future research into more principled and generalizable solutions
 Our main contributions are as follows:
-We propose the Unreasonable Math Problems (UMP) benchmark by minimally editing standard math datasets to create diverse and comprehensive unreasonable problems.
-We conduct comprehensive experiments and multi-dimensional analyses to examine how different types of LLMs behave when presented with unreasonable questions.
-We identify distinct failure modes and reveal that models with strong performance on well-posed tasks often behave unstably or fail to recognize unreasonableness when assumptions are flawed.
-We explore prompting and fine-tuning as strategies to investigate the potential of LLMs to detect unreasonableness.
+We propose the Unreasonable Math Problems (UMP) benchmark to more accurately and comprehensively evaluate how LLMs respond to mathematically unreasonable problem statements.
+We find that even high-performing models often fail to detect unreasonableness, or produce overconfident and overly verbose responses to unreasonable questions.
+We show that simple prompting or fine-tuning can partially mitigate these issues, but often introduce trade-offs such as decreased accuracy on well-posed problems, posing new challenges for future research.
 2 Unreasonable Math Problems(UMP) Benchmark
-While large language models have demonstrated strong performance on standard mathematical reasoning benchmarks, they often produce inaccurate or confusing responses when presented with mathematically unreasonable problems, questions that contain flawed assumptions, undefined variables, or logical inconsistencies (as show in Figure 1).
+While large language models have demonstrated strong performance on standard mathematical reasoning benchmarks, they often produce inaccurate or confusing responses when presented with mathematically unreasonable problems, questions that contain flawed assumptions, undefined variables, or logical inconsistencies (as shown in Figure 1).
 To systematically evaluate model behavior under such conditions, we construct the Unreasonable Math Problems (UMP) benchmark, which consists of 1000+ unreasonable math problems, each paired with its corresponding original version, focusing on assessing LLMs’ ability to detect and respond to irrational inputs.
-2.1 Data Category
+2.1 Types of Unreasonableness
 We identify five prevalent types of mathematical unreasonableness commonly found in LLM failure cases: (1) undefined variables, (2) illogical scenarios, (3) incorrect assumptions, (4) misinterpretation of units, and (5) inconsistent conditions.
-Each instance in our benchmark is represented as a quintuple , consisting of the original question and its answer , the unreasonable variant with its assigned type , and an explanation describing why is unreasonable.
+Each instance in our benchmark is represented as , consisting of the original question and its answer , the unreasonable variant with its assigned type , and an explanation describing why is unreasonable.
 Detailed definitions and examples for each category are provided in Appendix A.
-2.1.1 Data Generation
-Our data generation process is inspired by MetaMath (Yu et al., 2023b), which leverages LLMs to produce problem variants under controlled transformations.
+2.2 LLM-Guided Construction of Unreasonable Variants
+Our data construction process is inspired by MetaMath (Yu et al., 2023b), which leverages LLMs to produce problem variants under controlled transformations.
 As illustrated in Figure 2, we begin with test set questions drawn from four widely-used math benchmarks: GSM8K Cobbe et al.
@@
 Only variants whose similarity exceeds a predefined threshold are retained for human verification.
-2.1.2 Data Checking
-We perform manual validation with the assistance of three annotators, each with at least undergraduate-level mathematical training.
-For each example, annotators are provided with the original problem, its unreasonable variant, and a corresponding explanation, and are asked to verify whether the variant introduces genuine mathematical unreasonableness while preserving surface similarity.
-Only examples that meet both criteria are retained in the final benchmark.
-Our transformation rules are deliberately designed to introduce flaws that are logically problematic.
-As such, we filter out cases that either fail to introduce meaningful irrationality or make the unreasonableness too explicit, for example, variants that contain phrases like “-3 people” or “2.5 people,” which violate commonsense constraints in an obvious way.
-In these cases, we revise the questions by modifying entities, adjusting numerical values, or refining the phrasing to ensure that the unreasonable elements remain logically embedded but superficially plausible.
-2.1.3 Data Composition
+2.3 Validation Checking
+We manually verify each example to ensure that the unreasonable variant introduces genuine mathematical unreasonableness while maintaining surface similarity with the original problem.
+Only examples satisfying both criteria are retained.
+Our transformation rules are carefully designed to embed logical flaws.
+We discard cases that either lack meaningful irrationality or make the flaw overly explicit (see the annotation protocol in Appendix H).
+In such cases, we revise the examples by adjusting entities, numerical values, or phrasing to ensure the unreasonableness remains logically subtle yet plausible.
+2.4 Benchmark Composition
 Table 1 summarizes the distribution of unreasonableness types across different source datasets in the UMP benchmark.
@@
 This distribution emerges from the model generation and human filtering process.
-In particular, these two types of flaws tended to produce more plausible and contextually coherent questions, making them more likely to be retained during human recheck.
-In comparison, AMC23 and AIME24 contain relatively few test questions (with AIME24 having only 30), which naturally limits the number of examples that can be derived from these datasets.
-Moreover, problems in these sources are often abstract and symbolic, making them less amenable to realistic and controlled perturbation.
-Rather than artificially forcing unreasonable variants in such cases, we adopt a conservative strategy and include only those examples where natural modifications can be applied without compromising the original structure.
-This selective inclusion ensures that the final benchmark maintains both diversity and quality across a range of mathematical domains.
+In particular, these two types of flaws tend to produce more plausible and contextually coherent questions, making them more likely to be retained during human recheck.
+Compared to GSM8K and MATH, the AMC23 and AIME24 datasets contain far fewer test questions (e.g., AIME24 has only 30), which inherently limits the number of examples we can derive from them.
+In addition, these problems are often highly abstract and symbolic, making it difficult to apply natural, controlled perturbations without compromising their integrity.
+As a result, we only include modified examples when the unreasonableness can be introduced in a plausible manner.
+This selective inclusion ensures that the final benchmark remains both diverse and faithful to the original problem distributions.
 | Dataset | UV | IS | IA | MU | IC | Total |
@@
 | Total | 8.1% | 4.0% | 38.4% | 8.5% | 40.9% | 1129 |
-3 Experiments
-3.1 Experimental Setup
+3 Evaluating LLMs on Unreasonable Math Problems
+3.1 Evaluation Setup
 We evaluate a diverse set of LLMs spanning three major categories.
@@
 (2024).
-Reasoning-enhanced models are designed to support complex multi-step reasoning and include DeepSeek-R1 and its distilled variants (e.g., DeepSeek-R1-Distill-Qwen-7B) DeepSeek-AI (2025), as well as QwQ-32B-Preview Team (2025) and Grok-3-Reasoning.
-All models are evaluated on both the original problems and their unreasonable variants.111Details of model versions and inference configurations are provided in Appendix C.
+Reasoning-enhanced models are designed to support complex multi-step reasoning and include DeepSeek-R1 and its distilled variants (e.g., DeepSeek-R1-Distill-Qwen-7B) (DeepSeek-AI, 2025), as well as QwQ-32B-Preview (Team, 2025) and Grok-3-Reasoning (xAI, 2024).
+All models are evaluated on both the original problems and their unreasonable variants.111Details of model versions and inference configurations are provided in Appendix D.
 3.2 Evaluation Metrics
@@
 In conventional settings, the goal is to assess whether the model arrives at the correct answer.
-However, for unreasonable questions, the objective is to evaluate whether the model can recognize the irrationality embedded in the problem and respond appropriately by rejecting, questioning, or flagging the input as flawed.
+However, for unreasonable questions, the objective is to evaluate whether the model can recognize the unreasonability embedded in the problem and respond appropriately by rejecting, questioning, or flagging the input as flawed.
 To address this, we introduce two complementary metrics:
@@
 It is defined as the proportion of corresponding unreasonable variants that are correctly recognized as unreasonable, among those for which the model answers the original (reasonable) version correctly.
-This design controls for a factor that in some cases, the irrationality in a problem only becomes apparent during intermediate reasoning steps.
-A model that lacks the necessary mathematical competence may never reach the point where the irrationality is revealed.
+This design accounts for cases where the flaw only becomes apparent during intermediate reasoning steps.
+A model that lacks the necessary mathematical competence may never reach the point where the unreasonability is revealed.
 By restricting the evaluation to questions the model can already solve in their original form, relative score isolates the model’s ability to detect unreasonableness from its general problem-solving skill.
@@
 | Llama-3.1-8B-Instruct | ||||||||||||
-| DeepSeek-v3 | ||||||||||||
-| GPT-4o | 0.610 | |||||||||||
+| DeepSeek-v3 | 0.671 | 0.442 | ||||||||||
+| GPT-4o | 0.657 | 0.610 | ||||||||||
 | Claude-3.5-Sonnet | ||||||||||||
@@
 | DeepSeek-R1 | ||||||||||||
-| Grok3-Reasoning | 0.903 | |||||||||||
+| Grok3-Reasoning | 0.884 | 0.927 | 0.918 | 0.903 | ||||||||
 | Math Models | ||||||||||||
-| Qwen2.5-Math-1.5B-Instruct | 0.376 | |||||||||||
+| Qwen2.5-Math-1.5B-Instruct | 0.356 | 0.416 | 0.376 | |||||||||
 | MetaMath-Mistral-7B | ||||||||||||
@@
 | NuminaMath-7B-CoT | ||||||||||||
-| Qwen2.5-Math-7B-Instruct | ||||||||||||
+| Qwen2.5-Math-7B-Instruct | 0.458 | |||||||||||
 3.2.1 Absolute Score
-To evaluate how well a model identifies the unreasonableness of a given question, we adopt the LLM-as-a-judge paradigm Zheng et al.
-(2023), using DeepSeek-V3 to assess the quality of generated responses.
-Instead of a binary correct/incorrect judgment, we use a three-level rubric to better capture the nuances of model behavior:
-A: Correctly identifies the main source of unreasonableness, provides a coherent explanation, and proposes a valid correction without introducing new errors.
-B: Partially recognizes the unreasonableness or provides vague justification, but the explanation contains gaps, circular logic, or minor inconsistencies.
-C: Fails to recognize the core flaw in the problem, misinterprets the reasoning, or introduces new contradictions in the explanation or correction.
+Following the LLM-as-a-judge framework Zheng et al.
+(2023), we use DeepSeek-V3 to label each answer as A (correctly identifies the main source of unreasonableness, explains it coherently, and proposes a valid fix without new errors), B (partially detects or vaguely justifies the flaw, with gaps, circular logic, or minor inconsistencies), or C (fails to spot the core flaw, misreads the reasoning, or introduces new contradictions).
 Each model response receives a label , and we define a soft scoring function to assign partial credit:
@@
 3.2.2 Relative Score
-While the absolute score provides a general measure of robustness, it does not account for the model’s underlying mathematical competence.
-In many cases, the flaw in an unreasonable question only becomes apparent during intermediate reasoning steps.
 If a model lacks the skill to solve the original problem, it may never reach the point where the flaw is revealed.
-To control for this confound, we introduce the relative score, which measures a model’s ability to detect unreasonableness on the subset of questions it has already solved correctly in their original form.
+To control this factor, we introduce the relative score, which measures a model’s ability to detect unreasonableness on the subset of questions it has already solved correctly in their original form.
 Formally, following Yang et al.
@@
 As defined in Section 3.2, these behavioral scores capture a model’s ability to detect and respond appropriately to flawed problem setups.
-General-purpose models.
-Models such as GPT-4o, Claude-3.5-Sonnet, and DeepSeek-v3 perform well on well-posed problems (e.g., GPT-4o: 0.863 Acc) and show moderate ability to detect unreasonableness (0.603 Abs).
-However, smaller models like Qwen2.5-3B and LLaMA-3.1-8B struggle with behavioral metrics, indicating difficulty in recognizing flawed logic.
-This suggests that robustness to irrationality is not a natural outcome of instruction tuning and may require additional improvement.
-Reasoning-enhanced models.
-These models achieve consistently high scores, with DeepSeek-R1 and grok-3-reasoning reaching 0.813 Abs and 0.903 Rel, respectively.
-Their structured reasoning and verification steps appear effective for detecting logical flaws, though often at the cost of longer, more verbose outputs (see Section 4.2).
-Math-specialized models.
-Despite strong accuracy on original problems (e.g., Qwen2.5-Math-7B: 0.894 Acc), this group generally underperforms on unreasonable problems(score mostly below 0.30), revealing a gap between reasoning proficiency and irrationality detection.
-MetaMath-Mistral-7B exemplifies this limitation, with only 0.139 on Rel.
-Summary.
-We observe a trade-off between task accuracy and robustness to unreasonableness.
-Reasoning-enhanced models excel at flaw detection, while math-specialized models, despite high precision on well-posed problems, often fail to recognize inconsistencies in adversarial variants.
+We categorize evaluated models into three groups: general-purpose, reasoning-enhanced, and math-specialized.
+General models such as GPT-4o, Claude-3.5-Sonnet, and DeepSeek-v3 achieve strong accuracy on well-posed problems (e.g., GPT-4o: 0.863 Acc) and moderate robustness to unreasonableness (0.603 Abs), while smaller models like Qwen2.5-3B and LLaMA-3.1-8B struggle across behavioral metrics.
+Reasoning models (e.g., DeepSeek-R1, grok-3-reasoning) perform best on UMP, reaching 0.813 Abs and 0.903 Rel, benefiting from structured verification steps, but often generating overly verbose responses (see Section 4.1).
+In contrast, math-specialized models (e.g., Qwen2.5-Math-7B, MetaMath-Mistral-7B) excel in original task accuracy (up to 0.894 Acc) but perform poorly on UMP (mostly below 0.30), indicating a gap between precise reasoning and unreasonability detection.
+These results highlight a trade-off between accuracy on well-posed problems and sensitivity to unreasonableness, with reasoning models offering better detection at the cost of increased verbosity.
 4 Analysis
-4.1 Does Solving the Original Help?
-We investigate whether a model’s ability to solve the original version of a problem affects its capacity to detect flaws in the corresponding unreasonable variants.
-Intuitively, some flaws such as contradictions in intermediate steps may only be visible after progressing through the correct reasoning path.
-This motivates a direct comparison: how often does a model correctly identify unreasonableness when it has solved the original question, versus when it has not?
-To measure this, we revisit the set from Section 3.2, containing original questions that the model solves correctly with high confidence.
-We define its complement as the set of questions the model fails to solve.
-For each group, we compute the average score over their unreasonable variants, using the same scoring function .
-We then define the robustness gap as:
-where and denote the average scores over and , respectively.
-A larger gap indicates that the model’s ability to detect unreasonableness depends more strongly on its success in solving the original problem.
 4.1 Reasoning Models Tend to Overthink on Unreasonable Problems
@@
 (2025).
-We quantify overthinking by counting occurrences of reflective phrases such as “rethink” or “I misunderstood,” which indicate mid-generation self-correction (see Appendix D for keyword details).
+We quantify overthinking by counting occurrences of reflective phrases such as “rethink” or “I misunderstood,” which indicate mid-generation self-correction (see Appendix F for keyword details).
 Although we evaluate reflection frequency across all model types, only reasoning-enhanced models exhibit consistent and substantial increases on unreasonable problems.
@@
 This drop is especially pronounced in reasoning-enhanced models such as R1-Distill-Qwen-32B, where entropy decreases by 0.1–0.15 on average.
-These collapsed outputs often stem from repeated, ineffective reasoning loops, which reduce informativeness and mask the model’s failure to recognize flawed assumptions.
-Examples of such behavior are presented in Appendix I.1.
+These semantic collapse often stems from repeated, ineffective reasoning loops, which reduce informativeness and mask the model’s failure to recognize flawed assumptions.
+Examples of such behavior are presented in Appendix L.1.
 Token-Level Repetition.
@@
 Interestingly, Qwen2.5-Math-7B-Instruct, a math-specialized model with strong performance on well-posed questions, also shows a marked rise in repetition, suggesting that semantic instability is not limited to reasoning-heavy models.
-Appendix I.2 provides detailed case studies of this failure pattern.
-5 Can LLMs Detect Irrationality?
+Appendix L.2 provides detailed case studies of this failure pattern.
+4.3 Summary of Findings
+Our behavioral analysis uncovers several key patterns.
+First, general-purpose models often proceed with unwarranted confidence, failing to recognize flawed assumptions—a phenomenon we term unconscious of unreasonableness.
+Second, reasoning-enhanced models are prone to overthinking, repeatedly revising their reasoning in response to irrational inputs, which leads to verbosity and incoherence.
+Third, both reasoning and math-specialized models exhibit semantic collapse, characterized by increased token repetition and reduced entropy under unreasonable conditions.
+Finally, despite strong accuracy on well-posed tasks, math-specialized models often fail to detect subtle logical flaws, highlighting a disconnect between mathematical proficiency and robustness to flawed inputs.
+5 Can LLMs Detect unreasonability?
 Through the experiments above, we observe that most models perform poorly at recognizing unreasonableness in mathematical problems.
 To explore possible ways to improve this behavior, we begin with a simple probing experiment to test whether models possess the basic ability to judge flawed inputs.
-Based on this, we further investigate two strategies, prompt-based intervention and supervised fine-tuning—to enhance the model’s capacity to detect irrationality.
+Based on this, we further investigate two strategies, prompt-based intervention and supervised fine-tuning to enhance the model’s capacity to detect unreasonability.
 5.1 Probing Study
@@
 5.3 Fine-Tuning Based Method
-We further investigate whether the model’s ability to detect irrationality can be enhanced through supervised fine-tuning.
+We further investigate whether the model’s ability to detect unreasonability can be enhanced through supervised fine-tuning.
 To construct the training data, we pair unreasonable questions from GSM8K with carefully edited responses based on DeepSeek-R1 outputs.
@@
 This suggests a trade-off between flaw sensitivity and general problem-solving ability.
-Further training configurations and examples are detailed in Appendix F.
+Further training details are shown in Appendix I.
 Discussion.
@@
 The limitations of these methods underscore the inherent difficulty of the UMP benchmark and point to the need for more principled, generalizable flaw detection mechanisms that can integrate critical reasoning while preserving performance on standard tasks.
-6 Related Work
-6.1 LLMs as Math Problem Solvers
+6 Conclusion
+In this work, we investigate how LLMs behave when confronted with unreasonable mathematical problems.
+To facilitate this, we construct a benchmark comprising over 1,000 questions containing hidden unreasonability.
+Experimental results reveal that most models struggle to identify such unreasonableness.
+While reasoning-enhanced models are more likely to uncover hidden inconsistencies during multi-step reasoning, they often fall into repetitive reflection and overthinking behavior that ultimately hinders clarity and usefulness.
+To further explore whether models possess latent capabilities for detecting flaws, we experiment with critical-thinking prompts and supervised fine-tuning.
+Our findings suggest that many models do have the potential to detect unreasonable content, but this ability requires explicit activation.
+We hope that our benchmark serves as a valuable tool for evaluating both the trustworthiness and behavioral robustness of LLMs in the presence of unreasonable inputs.
+Limitations
+This study has several limitations.
+First, we do not include proprietary reasoning models such as those from OpenAI in our analysis of reasoning-enhanced systems.
+Since these APIs do not expose intermediate reasoning chains, comparing their behavior to open-source models that explicitly generate multi-step reasoning would introduce unfairness in both behavioral and metric-based evaluations.
+Second, due to hardware constraints, our intervention experiments such as prompt engineering and supervised fine-tuning are conducted only on 7B-scale models.
+While these experiments provide encouraging insights, scaling these methods to larger models and exploring more effective solutions for mitigating failure behaviors remain important directions for future work.
+Ethical Considerations
+Our paper explores how LLMs perform in the face of unreasonable mathematical problems.
+Sometimes the model may not realize the unreasonability in the math problem, which may result in incorrect answers that mislead users.
+Appendix A Category of UMP Benchmark
+Below is the description of the five categories.
+Each data point consists of an original question paired with its corresponding answer , a designated type of unreasonableness , a rephrased unreasonable version , and an explanation .
+Consequently, our benchmark comprises quintuples of the form .
+Appendix B Related Work
+B.1 LLMs as Math Problem Solvers
 Large language models have demonstrated impressive capabilities in solving math word problems and symbolic reasoning tasks.
@@
 These models outperform general-purpose LLMs on math benchmarks.
-6.2 Improve Models’ Inference Ability
+B.2 Improve Models’ Inference Ability
 Recent efforts have focused on enhancing the reasoning capabilities of large language models by structuring and extending their inference chains.
@@
 In addition, verifier-based strategies such as Outcome-Supervised Learning (Yu et al., 2023a) and post-hoc reflection mechanisms have been introduced to scrutinize intermediate steps and promote robust decision-making, especially when the model’s output chain may contain errors.
-6.3 Model Refusal Phenomenon
-The phenomenon of model refusal, where LLMs decline to answer certain prompts is central to ensuring AI safety and reliability.
-Refusals may stem from safety training, uncertainty or knowledge gaps or defects in the prompt itself.
+B.3 Model Refusal Phenomenon
+The phenomenon where LLMs decline to answer certain prompts is central to ensuring AI safety and reliability.
+Refusals may stem from safety training, uncertainty, knowledge gaps or defects in the prompt itself.
 Recent studies have explored both linear and nonlinear mechanisms behind refusal behaviors, including "refusal directions" in activation space and the use of "refusal tokens" to control response behavior at inference.
-However, over-refusal—where benign prompts are incorrectly rejected remains a key challengePasch (2025); Hildebrandt et al.
+However, over-refusal-where benign prompts are incorrectly rejected- remains a key challengePasch (2025); Hildebrandt et al.
 (2025).
@@
 (2025).
-7 Conclusion
-In this work, we investigate how large language models behave when confronted with unreasonable mathematical problems.
-To facilitate this, we construct a benchmark comprising over 1,000 questions containing hidden irrationality.
-Experimental results reveal that most models struggle to identify such unreasonableness.
-While reasoning-enhanced models are more likely to uncover hidden inconsistencies during multi-step reasoning, they often fall into repetitive reflection and overthinking behavior that ultimately hinders clarity and usefulness.
-To further explore whether models possess latent capabilities for detecting flaws, we experiment with critical-thinking prompts and supervised fine-tuning.
-Our findings suggest that many models do have the potential to detect unreasonable content, but this ability requires explicit activation.
-We hope that our benchmark serves as a valuable tool for evaluating both the reliability and behavioral robustness of LLMs in the presence of unreasonable inputs.
-Limitations
-This study has several limitations.
-First, we do not include proprietary reasoning models such as those from OpenAI in our analysis of reasoning-enhanced systems.
-Since these APIs do not expose intermediate reasoning chains, comparing their behavior to open-source models that explicitly generate multi-step thoughts would introduce unfairness in both behavioral and metric-based evaluations.
-Second, due to hardware constraints, our intervention experiments such as prompt engineering and supervised fine-tuning are conducted only on 7B-scale models.
-While these experiments provide encouraging insights, scaling these methods to larger models and exploring more effective solutions for mitigating failure behaviors remain important directions for future work.
-Ethical Considerations
-Our paper explores how llms perform in the face of unreasonable mathematical problems.
-Sometimes the model may not realize the irrationality in the math problem, so it may produce wrong answers and mislead the user.
-Appendix A Category of UMP Benchmark
-Below is the description of the five categories.
-Each of our data pieces is made up of five parts.
-An original question paired with its corresponding answer , along with a designated type of unreasonableness we aim to investigate, we systematically construct a rephrased, unreasonable question and provide an accompanying explanation .
-Consequently, our benchmark comprises quintuples of the form .
-Appendix B Probabilistic View from Prior Conditioning
+Appendix C Probabilistic View from Prior Conditioning
 We frame the model’s ability to recognize unreasonable problems as a conditional behavior dependent on its understanding of the original question.
@@
 It provides a graded measure of how reliably a model detects semantic flaws, assuming sufficient prior understanding of the original problem.
-Appendix C Model Versions and Inference Settings
+Appendix D Model Versions and Inference Settings
 We include both API-access and open-source models in our evaluation.
@@
 However, for reasoning-enhanced models, we observe that greedy decoding may result in unstable behaviors, such as excessively long outputs, especially in unreasonable problems.
-Based on these findings, we set temperature = 0.6 for all reasoning models during evaluation.(We provide further analysis in Appendix H.) We set , which means that the model must get all three responses right.
+Based on these findings, we set temperature = 0.6 for all reasoning models during evaluation.(We provide further analysis in Appendix K.) We set , which means that the model must get all three responses right.
 For API-based models, we maintain their default decoding settings.
 Due to access limitations and cost constraints, all the responses are generated in a single pass without sampling-based aggregation.
-Appendix D Keywords for Reflection Detection
+Appendix E Does Solving the Original Help?
+We investigate whether a model’s ability to solve the original version of a problem affects its capacity to detect flaws in the corresponding unreasonable variants.
+Intuitively, some flaws such as contradictions in intermediate steps may only be visible after progressing through the correct reasoning path.
+This motivates a direct comparison: how often does a model correctly identify unreasonableness when it has solved the original question, versus when it has not?
+To measure this, we revisit the set from Section 3.2, containing original questions that the model solves correctly with high confidence.
+We define its complement as the set of questions the model fails to solve.
+For each group, we compute the average score over their unreasonable variants, using the same scoring function .
+We then define the robustness gap as:
+where and denote the average scores over and , respectively.
+A larger gap indicates that the model’s ability to detect unreasonableness depends more strongly on its success in solving the original problem.
+Appendix F Keywords for Reflection Detection
 To quantify reflective behavior in model responses, we apply a keyword-based heuristic to detect phrases associated with self-correction, doubt, or reevaluation.
@@
 This keyword set was manually curated to capture a wide range of reflective behaviors and was applied in our analysis of model overthinking.
-Appendix E Similarity between Original Problems and Unreasonable Problems
+Appendix G Similarity between Original Problems and Unreasonable Problems
 To ensure that our constructed unreasonable problems remain as close as possible to their original counterparts, we conducted a similarity analysis between each pair of , where is the original question and is its unreasonable variant.
@@
 | Claude-3.5 | 0.503 | 0.305 | 0.198 |
-Appendix F Experiment Setup for Supervised Fine-Tuning
+Appendix H Annotation Protocol and Agreement
+To ensure the quality of the benchmark, three reviewers with mathematical training participated in a multi-stage validation process.
+The following criteria were jointly defined and applied during annotation:
+1.
+Remove examples that remain mathematically reasonable after rewriting but are mistakenly flagged as unreasonable by the model.
+2.
+Remove examples where the unreasonable elements are too obvious or violate basic commonsense (e.g., “2.33 people”).
+3.
+Remove examples where the rewritten version differs too much from the original, failing to preserve surface similarity.
+4.
+Revise examples with unclear or poorly explained unreasonable elements to make the flaw logically coherent and minimally invasive.
+An example is retained only if at least two out of three reviewers judged it as valid under the above criteria.
+Appendix I Experiment Setup for Supervised Fine-Tuning
 For the fine-tuning experiments, we use Qwen2.5-7B-Instruct and perform full-parameter supervised fine-tuning.
@@
 The well-posed samples are similarly constructed to ensure balanced learning and to mitigate overgeneralization of unreasonable detection behavior.
-Appendix G Template for Critical Prompting
+Appendix J Template for Critical Prompting
 Here we show our critical template for solving math problems.
-Appendix H Effect of Temperature on Model Behavior
+Appendix K Effect of Temperature on Model Behavior
 We analyze how different decoding temperatures affect model behavior, focusing on two representative models: Qwen2.5-7B-Instruct (general-purpose) and Qwen-Distill-7B (reasoning-enhanced).
@@
 Overall, a temperature of 0.6 yields the most balanced performance across models, aligning with the default setting recommended in the DeepSeek official recommendation.
-Appendix I Case Study
-I.1 An Example of Overthinking
+Appendix L Case Study
+L.1 Examples of Overthinking
 Table 6 and Table 7 presents representative case of DeepSeek-R1’s behaviors when confronted with unreasonable problems.
@@
 Such repetitive reasoning over a trivial computation is clearly unnecessary and results in substantial computational overhead.
-I.2 The unstable of math-specified model
+L.2 The Unstability of Math-Specified Model
 Table 8 illustrates the behavior of Qwen2.5-Math-7B-Instruct when confronted with unreasonable math problems.
 As shown, the model often becomes paralyzed in the presence of flawed inputs, failing to identify the source of unreasonableness and unable to proceed with a meaningful response.
-I.3 Model Unconsciousness of Problem Errors
-Tables 9, 10, present examples where the model completely fails to recognize the errors embedded in the questions.
+L.3 Model Unconsciousness of Problem Errors
+Tables 9 and 10, present examples where the model completely fails to recognize the errors embedded in the questions.
 These cases suggest that the model entirely overlooks the unreasonableness of the input and proceeds as if the problems were well-posed.
@@
 They couldn’t eat more than 4, so all pieces are gone.**Final Answer: 0** |
-| Column 1 | Column 2 |
 | Unreasonable Problem | Jenna and her mother picked some apples from their apple farm. Jenna picked half as many apples as her mom. If her mom got 15 apples, how many apples did they both pick? |
@@
 |
-| Column 1 | Column 2 |
 | Unreasonable Problem | Compute in your head. |
@@
 Yeah, I think that’s it.**Final Answer** |
-| Column 1 | Column 2 |
 | Unreasonable Problem | Two candidates are running for class representative at Sarai’s school.
@@
 |
-| Column 1 | Column 2 |
 | Unreasonable Problem | A regular hexagon can be divided into six equilateral triangles.
@@
 |
-| Column 1 | Column 2 |
 | Unreasonable Problem | What is the least positive integer multiple of 30 that can be written with only the digit 2? |
@@
 |
-| Column 1 | Column 2 |
 | Unreasonable Problem | John has 3 boxes. Each box is 5 grams by 6 grams by 4 grams. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? |
@@
 |
-| Column 1 | Column 2 |
 | Unreasonable Problem | Cars have lined up on the motorway.
@@
 |
-| Column 1 | Column 2 |
 | Unreasonable Problem | A salesman bought a case of 48 sneakers for $576.
@@
 Final Answer: |
-| Column 1 | Column 2 |
 | Unreasonable Problem | Four books are arranged on a shelf.
@@
 - Total Thickness = 3100 mm + 50 mm + 45 mm + 6200 mm = mm" |
-| Column 1 | Column 2 |
 | Unreasonable Problem | Uncle Ben has four horses.
@@
 Determine the number of bags needed: - Weight of one bag: 50 pounds - Number of bags needed: Final Answer: |
-Appendix J Example output by different LLM
-In this section, we will demonstrate how various models perform on unreasonable problems using the conventional Chain of Thought approach compared to adding a critical instruction method.
-Tables 16 display the outputs from Qwen2.5-7B-Instruct using two different prompts.
+Appendix M Example output by different LLM
+In this section, we demonstrate how various models perform on unreasonable problems using the conventional Chain of Thought approach compared to adding a critical instruction method.
+Table 16 displays the outputs from Qwen2.5-7B-Instruct using two different prompts.
 It is evident that the outputs generated by adding critical instruction exhibit marked enhancements.
@@
 |
-Appendix K Evaluation Prompt Details
+Appendix N Evaluation Prompt Details
 We use an LLM-as-a-judge setup to assess whether a model response reasonably identifies the unreasonableness in the modified question.
@@
 For each dataset, we design corresponding in-context examples to better assess the model’s performance.
-Appendix L Prompt for Unreasonable Problem Generation
+Appendix O Prompt for Unreasonable Problem Generation
 We present the detailed prompt used for generating unreasonable math problems via LLMs.
 Note that the actual prompt is written in Chinese, and the English version shown below is a direct translation for presentation purposes.
-Appendix M Examples in UMP benchmark
+Appendix P Examples in UMP benchmark
 In this section we will show one unreasonable problem for each category in our benchmark in table 17.
@@
 Our intention is that this explanation aids test creators in reviewing and refining their problems to ensure logical consistency.
-Appendix N Query format for Pre-validation experiment
+Appendix Q Query format for Pre-validation experiment
 Our prompt is as follows: Please decide whether the following statement is reasonable:{d}.Where “d” is the description shown in the following table